### v1

# SymplecticDuality.bib

Abstract.
We review the notion of symplectic duality earlier introduced in the context of topological recursion. We show that the transformation of symplectic duality can be expressed as a composition of $x-y$ dualities in a broader context of log topological recursion. As a corollary, we establish nice properties of symplectic duality: various convenient explicit formulas, invertibility, group property, compatibility with topological recursion and KP integrability. As an application of these properties, we get a new and uniform proof of topological recursion for large families of weighted double Hurwitz numbers; this encompasses and significantly extends all previously known results on this matter.
1. Introduction
Topological recursion (TR) is a procedure that associates to a so-called spectral curve data that consists of a Riemann surface $\Sigma$ and a pair of meromorphic functions $(x,y)$ a system of symmetric meromorphic differentials [EO]. It was originally developed as a computational tool for the cumulants of matrix models, but since then it evolved to a new and active area of research by itself that provides a universal interface that relates a huge class of algebraic and combinatorial enumerative problems to the intersection theory of moduli spaces and integrable systems.
In the last few years the structural development of the theory of topological recursion turned it into a powerful computational tool that not just extends the variety of problems governed by TR but also provides explicit closed formulas for the $n$-point generating functions (as opposed to recursive computation that TR is meant to give) and closed algebraic relations among them. Among the most important developments are the following ones:
$x-y$ duality:
The $x-y$ duality traces what happens with the correlation differentials under the swap of $x$ and $y$. An explicit closed formula was conjectured in [borot2023functional], and it was soon proved in [hock2022xy] in a special case, substantially simplified in [hock2022simple], and proved in general case in [ABDKS1]. This formula gained further reformulations and applications (in particular, to KP integrability) in [hock2023laplace, ABDKS3].
symplectic duality:
Symplectic duality was introduced in [BDKS4] as a generalization of the $x-y$ duality that depends on an additional rational function. It was used mostly in the context of vacuum expectation values to prove topological recursion for weighted double Hurwitz numbers / weighted maps and constellations with internal faces (see also [bonzom2022topological], though the latter reference does not cover the most general cases like the $r$-spin Hurwitz numbers), and for the so-called generalized fully simple maps in [ABDKS2].
logarithmic TR:
It was first observed in [hock2023xy] and further structurally developed in [ABDKS4] that one can extend the realm of applications of $x-y$ duality in topological recursion beyond the case of meromorphic $x$ and $y$. It is sufficient to assume that $dx$ and $dy$ are meromorphic, thus $x$ and $y$ are allowed to have logarithmic singularities. In this case the formulation of topological recursion needs an adjustment which we call the log topological recursion.
The main goal of this paper is to show that more general symplectic duality can, in fact, be represented as a sequence of $x-y$ dualities, with eventual corrections coming from the necessity to use logarithmic topological recursion. This unifies the existing approaches to explicit closed formulas for the $n$-point differentials, identifies competing closed formulas provided in the literature (which is especially remarkable since these formulas naturally occur in quite different contexts), upgrades symplectic duality to a group action on the spectral curve data and associated differentials (and brings it far beyond the original setup of vacuum expectation values where it was first developed), and shows that the symplectic duality preserves the KP integrability. The results of the present paper also serve as another argument in favor of the claim that logarithmic TR as defined in [ABDKS4] is indeed the natural extension of TR to spectral curves with logarithmic singularities.
As an application of our results on symplectic duality we get a new and uniform proof of (logarithmic) topological recursion for large families of weighted double Hurwitz numbers. This encompasses all previously known results on this matter (including the families studied in [BDKS2]; also see [BDKS2, Section 5.2] in particular for the special cases of those families studied in the prior literature), but also quite significantly extends beyond that, as we can, in particular, now remove various technical restrictions imposed before. This proof is also quite a bit nicer than everything available before, since even in [BDKS2] one had to carefully analyze various possible poles (separately for different families) to prove the projection property, while now we give a much simpler and more general and more conceptual proof.
We also give a new way how to identify various formulas for Kontsevich–Witten $n$-point functions, and we give a new proof of a recent theorem of Bouchard, Kramer, and Weller [BKW] that transalgebraic topological recursion reproduces in the limit the so-called Atlantes Hurwitz numbers (which do not satisfy the original TR by themselves), introduced in [ALS].
1.1. Organization of the paper
In Section 2 we remind the definitions of the ordinary and logarithmic topological recursion. In Section 3 we provide a review of the main results on the $x-y$ duality. In Section 4 we recall the notion of symplectic duality. In Section 5 we state and prove the main result (Theorem 5.1) of this paper that reduces symplectic duality to $x-y$ duality. In Section 6 we append the results of the previous section with an analysis of the (logarithmic) projection property in order to establish the connection to (log) topological recursion; as an application we get a new proof of (log) topological recursion for various general families of weighted double Hurwitz numbers. Finally, in Section 7 we collect some further examples.
1.2. Notation
Throughout the text we use the following notation:
  • $\llbracket{n}\rrbracket$ denotes $\{1,\dots,n\}$.
  • $z_{I}$ denotes $\{z_{i}\}_{i\in I}$ for $I\subseteq\llbracket{n}\rrbracket$.
  • $[u^{d}]$ denotes the operator that extracts the corresponding coefficient from the whole expression to the right of it, that is, $[u^{d}]\sum_{i=-\infty}^{\infty}a_{i}u^{i}\coloneqq a_{d}$.
  • $\mathop{\big{\lfloor}_{{u}\to{v}}}$ denotes the operator of substitution (or restriction), that is, $\mathop{\big{\lfloor}_{{u}\to{v}}}f(u)\coloneqq f(v)$.
  • $\mathcal{S}(u)$ denotes $u^{-1}(e^{u/2}-e^{-u/2})$.
  • $d_{i}$ stands for taking the differential with respect to the $i$-th variable.
1.3. Acknowledgments
A. A. was supported by the Institute for Basic Science (IBS-R003-D1). B.B. was supported by the ISF Grant 876/20. B.B., P.D.-B., and M.K. were supported by the Russian Science Foundation (grant No. 24-11-00366). S. S. was supported by the Netherlands Organization for Scientific Research.
We thanks V. Bouchard, A. Hock, R. Kramer, and Q. Weller for useful discussions.
2. Topological recursion and logarithmic topological recursion
2.1. Topological recursion
Let us recall the definition of topological recursion (TR) in terms of the loop equations and projection property. We follow the exposition in [ABDKS1, Section 5].
Let $x$ and $y$ be meromorphic functions on a compact Riemann surface $\Sigma$, and $\{\omega^{(g)}_{m}\}$, $g\geq 0$, $m\geq 1$, $02g-2+m>$ 0 be a given system of symmetric meromorphic differentials ($\omega^{(g)}_{m}$ is defined on $\Sigma^{m}$). We also define $\omega^{(0)}_{1}\coloneqq ydx$ and $\omega^{(0)}_{2}\coloneqq B$, where $B$ is the so-called Bergman kernel, that is, the unique symmetric meromorphic bi-differential on $\Sigma^{2}$ with a double pole along the diagonal with bi-residue $1$ and no further singularities, whose $\mathfrak{A}$-periods are all equal to 00.
Remark 2.1.
Note that we use the definition $\omega^{(0)}_{1}\coloneqq ydx$ in the present paper, as opposed to a different convention $\omega^{(0)}_{1}\coloneqq-ydx$ used, e.g., in [ABDKS1, ABDKS4].
Both conventions are widely used in the literature.
Assume that all zeros of $dx$ are simple and $y$ is regular at the zeros of $dx$ and the zeros of $dx$ and $dy$ are disjoint. Denote the zeros of $dx$ by $p_{1},\dots,p_{N}\in\Sigma$ and let $\sigma_{i}$ be the deck transformations of $x$ near $p_{i}$.
Definition 2.2.
We say that the system of symmetric meromorphic differentials $\{\omega^{g}_{m}\}$, $g\geq 0$, $m\geq 1$ satisfies
  • the linear loop equations, if for any $g,m\geq 0$, $i=1,\dots,N$
    (1) $\omega^{(g)}_{m+1}(z_{\llbracket m\rrbracket},z)+\omega^{(g)}_{m+1}(z_{\llbracket m\rrbracket},\sigma_{i}(z))$
    is holomorphic at $z\to p_{i}$ and has at least a simple zero in $z$ at $z=p_{i}$;
  • the quadratic loop equations, if for any $g,m\geq 0$, $i=1,\dots,N$, the quadratic differential in $z$
    (2) $\omega^{(g-1)}_{m+2}(z_{\llbracket m\rrbracket},z,\sigma_{i}(z))+\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ I_{1}\sqcup I_{2}=\llbracket m\rrbracket\end{subarray}}\omega^{(g_{1})}_{|I_{1}|+1}(z_{I_{1}},z)\omega^{(g_{2})}_{|I_{2}|+1}(z_{I_{2}},\sigma_{i}(z))$
    is holomorphic at $z\to p_{i}$ and has at least a double zero at $z=p_{i}$.
Remark 2.3.
Note that if it is already known that a given system of symmetric meromorphic differentials $\{\omega^{(g)}_{m}\}$, $g\geq 0$, $m\geq 1$ satisfies the linear loop equations, then the quadratic loop equations are equivalent to the following statement (see [DKPS-rspin]): for any $g,m\geq 0$, $i=1,\dots,N$, the expression
(3) $\mathop{\big{\lfloor}_{{w}\to{z}}}\Delta_{i,z}\Delta_{i,w}\left(\omega^{(g-1)}_{m+2}(z_{\llbracket m\rrbracket},z,w)+\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ I_{1}\sqcup I_{2}=\llbracket m\rrbracket\end{subarray}}\omega^{(g_{1})}_{|I_{1}|+1}(z_{I_{1}},z)\omega^{(g_{2})}_{|I_{2}|+1}(z_{I_{2}},w)\right)$
is holomorphic at $z\to p_{i}$ and has at least a double zero at $z=p_{i}$. Here $\Delta_{i,z}f(z):=f(z)-f(\sigma_{i}(z))$.
Definition 2.4.
We say that $\{\omega^{(g)}_{m}\}$ satisfies the blobbed topological recursion [BS-blobbed], if it satisfies the linear and quadratic loop equations.
Under the assumption that all zeros $\{p_{1},\dots,p_{N}\}$ of $dx$ are simple, the statement that $\{\omega^{(g)}_{m}\}$ satisfies the blobbed topological recursion implies that the principal parts near the zero locus of $dx$ of each $\omega^{(g)}_{m}$, $02g-2+m>$ 0, in each variable are uniquely determined by $\omega^{(g^{\prime})}_{m^{\prime}}$ with $2g^{\prime}-2+m^{\prime}<2g-2+m$.
Definition 2.5.
We say that $\{\omega^{(g)}_{m}\}$ satisfy the original topological recursion (TR) on the spectral curve $(\Sigma,x,y)$, if in addition to the blobbed topological recursion they satisfy the so-called projection property: for any $g\geq 0$, $m\geq 1$, $02g-2+m>$ 0
(4) $\omega^{(g)}_{m}(z_{\llbracket{m}\rrbracket})=\sum_{i_{1},\dots,i_{m}=1}^{N}\Bigg{(}\prod_{j=1}^{m}\mathop{\rm res}_{z_{j}^{\prime}\to p_{i_{j}}}\int^{z_{j}^{\prime}}_{p_{i_{j}}}B(\cdot,z_{j})\Bigg{)}\omega^{(g)}_{m}(z^{\prime}_{\llbracket{m}\rrbracket}).$
In other words, we say that in the case of original topological recursion the meromorphic differentials $\omega^{(g)}_{m}$, $02g-2+n>$ 0, have no other poles than at $\{p_{1},\dots,p_{N}\}$, and all their $\mathfrak{A}$-periods are equal to zero.
2.2. Logarithmic topological recursion
2.2.1. Motivation of log topological recursion
Topological recursion relations depend on the differentials $dx$, $dy$ only, but not the functions $x$ and $y$ themselves, and they are well defined if $dx$ and $dy$ are global meromorphic differentials with possibly nonzero residues. In the latter case the functions $x$ and $y$ are not univalued and posses logarithmic singularities. In the presence of this sort of singularities it is useful to consider, along with the original topological recursion, its variation that we call logarithmic topological recursion (LogTR), introduced in [ABDKS4] and motivated by the results of [hock2023xy]. It turns out that LogTR (which coincides with the original TR for meromorphic $x$ and $y$ but differs if $y$ has logarithmic singularities of certain type described below) is the proper version of topological recursion if one wants to consider the $x-y$ duality or the symplectic duality which are discussed below. That is, as proved in [ABDKS4], the $x-y$ duality formula holds for LogTR in the presence of logarithmic singularities in the same form as for meromorphic $x$ and $y$. In the present paper we prove a similar statement regarding the more general symplectic duality. Since these statements have numerous implications for many enumerative problems, as discussed in [hock2023xy, ABDKS4] and the present paper respectively, one might argue that LogTR is the correct definition in the presence of logarithmic singularities.
2.2.2. Definition of log topological recursion
Assume that $dx$ and $dy$ are globally defined meromorphic $1$-forms on a compact Riemann surface $\Sigma$ with possibly non-vanishing residues at some points. We still assume that all zeros $\{p_{1},\dots,p_{N}\}$ of $dx$ are simple, $dy$ is regular at the zero locus of $dx$ and the zero loci of $dx$ and $dy$ are disjoint.
Definition 2.6.
We say that the primitive $y$ of the differential $dy$ posses logarithmic singularity at some point $a$ on $\Sigma$ if $dy$ has a pole at $a$ with nonzero residue. A logarithmic singularity of $y$ is called LogTR-vital if this pole of $dy$ is simple and $dx$ has no pole at this point.
Let $a_{1},\dots,a_{M}$ be the LogTR-vital singular points of $y$. We denote the residues of $dy$ at these points by $\alpha_{1}^{-1},\dots,\alpha_{M}^{-1}$, respectively. That is, the principal part of $dy$ near $a_{i}$ is given by $\alpha_{i}^{-1}dz/(z-a_{i})$ in any local coordinate $z$.
Consider the germs of the principal parts of meromorphic $1$-forms in the neighborhoods of $a_{i}$, $i=1,\dots,M,$ defined as the principal parts of coefficients of positive powers of $\hbar$ in the following expressions:
(5) $\left(\frac{1}{\alpha_{i}\mathcal{S}(\alpha_{i}\hbar\partial_{x})}\log(z-a_{i})\right)dx.$
Definition 2.7.
Let $\{{}^{\log{}}\omega^{(g)}_{m}\}$, $g\geq 0$, $m\geq 1$, $02g-2+m>$ 0 be a given system of symmetric meromorphic differentials, and set ${}^{\log{}}\omega^{(0)}_{1}=ydx$, ${}^{\log{}}\omega^{(0)}_{2}=B$. We say that $\{{}^{\log{}}\omega^{(g)}_{m}\}$ satisfies the logarithmic topological recursion (LogTR) on the spectral curve $(\Sigma,x,y)$, if $\{{}^{\log{}}\omega^{(g)}_{m}\}$ satisfy the blobbed topological recursion and additionally they satisfy the so-called logarithmic projection property: for any $g\geq 0$, $m\geq 1$, $02g-2+m>$ 0
(6) ${}^{\log{}}\omega^{(g)}_{m}(z_{\llbracket{m}\rrbracket})$ $=\sum_{i_{1},\dots,i_{m}=1}^{N}\Bigg{(}\prod_{j=1}^{m}\mathop{\rm res}_{z_{j}^{\prime}=p_{i_{j}}}\int^{z_{j}^{\prime}}_{p_{i_{j}}}B(\cdot,z_{j})\Bigg{)}{}^{\log{}}\omega^{(g)}_{m}(z^{\prime}_{\llbracket{m}\rrbracket})$
$\quad+\delta_{m,1}[\hbar^{2g}]\sum_{i=1}^{M}\mathop{\rm res}_{z^{\prime}=a_{i}}\Bigg{(}\int^{z^{\prime}}_{a_{i}}B(\cdot,z_{1})\Bigg{)}\left(\frac{1}{\alpha_{i}\mathcal{S}(\alpha_{i}\hbar\partial_{x^{\prime}})}\log(z^{\prime}-a_{i})\right)dx^{\prime}.$
In other words, we say that in the case of logarithmic topological recursion the meromorphic differentials ${}^{\log{}}\omega^{(g)}_{m}$, $02g-2+m>$ 0, $m\geq 2$, have no other poles than at $\{p_{1},\dots,p_{N}\}$ and all their $\mathfrak{A}$-periods are equal to zero. For $g\geq 1$ and $m=1$ we have a bit different setup: they have poles at $\{p_{1},\dots,p_{N}\}$, and also at the points $\{a_{1},\dots,a_{M}\}$. The principal parts at the latter points are equal to the principal parts of expressions given by (5), and all their $\mathfrak{A}$-periods are equal to zero.
Remark 2.8.
Note that if $a$ is a simple pole of $dy$ and $dx$ has also a pole at this point, that is, singularity of $y$ is not LogTR-vital, then the principal part of $[\hbar^{2g}]\bigl{(}\frac{1}{\alpha\mathcal{S}(\alpha\hbar\partial_{x})}\log(z-a)\bigr{)}dx$ with $g>0$ is equal to zero, so we could formally include these points to the list of LogTR-vital singularities of $y$ without changing relation of recursion. But it is important to underline that such singular points do not contribute to the recursion and do not affect Equation (6).
Remark 2.9.
The poles of order greater than $1$ are excluded from the logarithmic projection property by similar local considerations. Indeed, assume that the meromorphic form $dy$ varies in a family such that two simple poles collapse together producing a pole of order $2$. Then, it is easy to see by local computations that the residues at these two poles tend to infinity when the points approach one another. It follows that the limiting contribution of these two points to (5) is trivial, independently of whether the limiting pole of order two has a residue or not. The same argument also works for the higher order poles.
3. Review of $x-y$ duality
In this section we remind the reader a definition and the main properties of the $x-y$ duality, see [ABDKS1, ABDKS3, ABDKS4] for details.
3.1. Spectral curve data
A spectral curve data is a compact Riemann surface (a smooth complex algebraic curve) $\Sigma$ equipped with a pair of functions $(x,y)$ on it such that $dx$ and $dy$ are meromorphic $1$-forms. We do not assume that $x$ and $y$ are themselves meromorphic; these functions could be multivalued and possess logarithmic singularities.
Let $\{\omega^{(g)}_{n}\}$ be a system of symmetric meromorphic $n$-differentials on $\Sigma^{n}$ defined for $g\geq 0$, $n\geq 1$, $(g,n)\neq(0,1)$ such that $\omega^{(g)}_{n}-\delta_{(g,n),(0,2)}\frac{dx_{1}dx_{2}}{(x_{1}-x_{2})^{2}}$ is regular on the diagonals in $\Sigma^{n}$. We extend this system of differentials to the case $(g,n)=(0,1)$ by setting
(7) $\omega^{(0)}_{1}=y\,dx$
even though this form may not be well defined as a meromorphic one. Let us stress here that we do not assume that the differentials satisfy any topological recursion.
Remark 3.1.
The definition of the spectral curve data that can be found in the literature often includes also a choice for the 2-differential $B=\omega^{(0)}_{2}$. In most cases of our particular interest, we have $\Sigma=\mathbb{C}P^{1}$ and $\omega^{(0)}_{2}$ is uniquely defined by $\omega^{(0)}_{2}=\frac{dz_{1}dz_{2}}{(z_{1}-z_{2})^{2}}$ for an arbitrary affine coordinate $z$. By that reason we will often denote by $z_{i}$ a point on the $i$th copy of the Cartesian product $\Sigma^{n}$ identifying this point with the value of the coordinate $z$ at this point, and denote $x_{i}=x(z_{i})$, $y_{i}=y(z_{i})$. Note, however, that the basic definitions and relations of this paper can be applied to a spectral curve of any genus (or to a different choice for $\omega^{(0)}_{2}$ in the case of a rational spectral curve).
3.2. Extended $n$-point differentials
The $x-y$ duality construction involves summation over hypergraphs as an intermediate step.
Definition 3.2.
Given a system of differentials $\{\omega^{(g)}_{n}\}$, the extended $n$-point differentials $\Omega_{n}=\sum_{g=0}^{\infty}\hbar^{2g-2+n}\Omega^{(g)}_{n}$, $\mathbb{W}_{n}=\sum_{g=0}^{\infty}\hbar^{2g-2+n}\mathbb{W}^{(g)}_{n}$ are defined by
(8) $\Omega_{n}(z^{+}_{\llbracket{n}\rrbracket},z^{-}_{\llbracket{n}\rrbracket})$ $=\prod_{i=1}^{n}\tfrac{\sqrt{dx_{i}^{+}dx_{i}^{-}}}{x_{i}^{+}-x_{i}^{-}}\;\sum_{\Gamma}\tfrac{1}{|\mathrm{Aut}(\Gamma)|}\prod_{e\in E(\Gamma)}\sum_{\tilde{g}=0}^{\infty}\hbar^{2\tilde{g}-2+|e|}\int\limits_{z_{e(1)}^{-}}^{z_{e(1)}^{+}}\dots\int\limits_{z_{e(|e|)}^{-}}^{z_{e(|e|)}^{+}}\tilde{\omega}^{(\tilde{g})}_{|e|},$
(9) $\mathbb{W}_{n}(z_{\llbracket{n}\rrbracket},u_{\llbracket{n}\rrbracket})$ $=\prod_{i=1}^{n}\Bigl{(}e^{u_{i}(\mathcal{S}(u_{i}\hbar\partial_{x_{i}})-1)y_{i}}\mathop{\big{\lfloor}_{{z_{i}^{\pm}}\to{e^{\pm\frac{u_{i}\hbar}{2}\partial_{x_{i}}}z_{i}}}}\Bigr{)}\;\Omega_{n}(z^{+}_{\llbracket{n}\rrbracket},z^{-}_{\llbracket{n}\rrbracket}).$
Here
  • $x_{i}=x(z_{i})$, $y_{i}=y(z_{i})$, where $z_{i}$ is a point on the $i$-th copy of $\Sigma$ in $\Sigma^{n}$, and similarly $x_{i}^{\pm}=x(z_{i}^{\pm})$, where $(z_{i}^{+},z_{i}^{-})$ are points near the diagonal on the $i$-th copy of $\Sigma^{2}$ in $(\Sigma^{2})^{n}$.
  • The sum is taken over all connected bipartite graphs (or hypergraphs or just graphs with multiedges, for brevity) $\Gamma$ with $n$ labeled vertices and multiedges of arbitrary index $\geq 1$, where the index of a multiedge is the number of its “legs” and we denote it by $|e|$.
  • For a multiedge $e$ with index $|e|$ we control its attachment to the vertices by the associated map $e\colon\llbracket{|e|}\rrbracket\to\llbracket{n}\rrbracket$ that we denote also by $e$, abusing notation (so $e(j)$ is the label of the vertex to which the $j$-th “leg” of the multiedge $e$ is attached).
  • Terms with $(g,n)=(0,1)$ are excluded from the summation, that is, we set $\tilde{\omega}^{(0)}_{1}=0$, and the meaning of $\tilde{\omega}^{(0)}_{|e|}$ in the case $|e|=2$ depends on whether the two legs $e(1)$ and $e(2)$ are the same or not.
  • We set $\tilde{\omega}^{(0)}_{|e|}(z_{1},z_{2})=\omega^{(0)}_{2}(z_{1},z_{2})-\frac{dx_{1}dx_{2}}{(x_{1}-x_{2})^{2}}$ if $e(1)=e(2)$, and we set $\tilde{\omega}^{(0)}_{|e|}=\omega^{(0)}_{2}$ otherwise. For all $(g,n)\neq(0,2)$ or $(0,1)$ we simply have $\tilde{\omega}^{(g)}_{n}=\omega^{(g)}_{n}$.
  • $|\mathrm{Aut}(\Gamma)|$ stands for the number of automorphisms of $\Gamma$.
Remark 3.3.
The arguments $(z^{+}_{i},z^{-}_{i})$ of $\Omega_{n}$ are treated as a point of the Cartesian square $\Sigma^{2}$ and are considered in a certain vicinity of the diagonal $z_{i}^{+}=z_{i}^{-}$. The arguments $(z_{i},u_{i})$ of $\mathbb{W}_{n}$ provide a particular parametrization of this vicinity through the substitutions
(10) $z_{i}^{+}=e^{\frac{u_{i}\hbar}{2}\partial_{x_{i}}}z_{i},\qquad z_{i}^{-}=e^{-\frac{u_{i}\hbar}{2}\partial_{x_{i}}}z_{i}.$
Remark 3.4.
The half-differentials $\sqrt{dx_{i}^{\pm}}$ entering $\Omega_{n}$ are understood formally. They attain a more definite meaning after the above substitution only. Namely, we have
(11) $\mathop{\big{\lfloor}_{{z_{i}^{\pm}}\to{e^{\pm\frac{u_{i}\hbar}{2}\partial_{x_{i}}}z_{i}}}}x_{i}^{\pm}=x_{i}\pm\frac{u_{i}\hbar}{2},\qquad\mathop{\big{\lfloor}_{{z_{i}^{\pm}}\to{e^{\pm\frac{u_{i}\hbar}{2}\partial_{x_{i}}}z_{i}}}}\frac{\sqrt{dx_{i}^{+}dx_{i}^{-}}}{x_{i}^{+}-x_{i}^{-}}=\frac{dx_{i}}{u_{i}\hbar}.$
Remark 3.5.
We regard $\mathbb{W}_{n}$’s as $n$-differentials on $\Sigma^{n}$ depending on additional parameters $u_{1},\dots,u_{n}$. One can see that they are expressed in terms of the differentials $\omega^{(\tilde{g})}_{\tilde{n}}$ and their derivatives (no actual integration is needed). This fact follows from the equality
(12) $\prod_{i=1}^{\tilde{n}}\mathop{\big{\lfloor}_{{\tilde{z}_{i}^{\pm}}\to{e^{\pm\frac{u_{e(i)}\hbar}{2}\partial_{x_{e(i)}}}z_{e(i)}}}}\;\int\limits_{\tilde{z}_{1}^{-}}^{\tilde{z}_{1}^{+}}\dots\int\limits_{\tilde{z}_{\tilde{n}}^{-}}^{\tilde{z}_{\tilde{n}}^{+}}\tilde{\omega}^{(\tilde{g})}_{\tilde{n}}=\prod_{i=1}^{\tilde{n}}\bigl{(}\mathop{\big{\lfloor}_{{\tilde{z}_{i}}\to{z_{e(i)}}}}u_{e(i)}\hbar\,\mathcal{S}(u_{e(i)}\hbar\,\partial_{\tilde{x}_{i}})\bigr{)}\;\frac{\tilde{\omega}^{(\tilde{g})}_{\tilde{n}}(\tilde{z}_{\llbracket{\tilde{n}}\rrbracket})}{\prod_{i=1}^{\tilde{n}}d\tilde{x}_{i}}.$
Note also that the coefficient of each power of $\hbar$ in $\mathbb{W}_{n}-\delta_{n,1}\frac{dx_{1}}{u_{1}\hbar}$ is a polynomial in $u$-variables, thus $\mathbb{W}^{(g)}_{n}$ for any $g$ and $n$ is a differential polynomial in $\omega$’s.
Remark 3.6.
A nice property of the (half-)differentials $\Omega_{n}$ is that they are actually independent of $x$ and $y$ and are uniquely determined by the differentials $\omega^{(g)}_{n}$ with $(g,n)\neq(0,1)$. Indeed, a different choice of the $x$-function changes both the factors $\frac{\sqrt{dx_{i}^{+}dx_{i}^{-}}}{x_{i}^{+}-x_{i}^{-}}$ and the regularization terms $\frac{d\tilde{x}_{1}d\tilde{x}_{2}}{(\tilde{x}_{1}-\tilde{x}_{2})^{2}}$ entering the definition of $\tilde{\omega}^{(0)}_{2}$. An easy computation shows that these two changes compensate one another, see [ABDKS3, Equation (3.10)].
Example 3.7.
Assume that $\omega^{(g)}_{n}=0$ for all $(g,n)$ except $\omega^{(0)}_{1}=ydx$ and $\omega^{(0)}_{2}=\frac{dz_{1}dz_{2}}{(z_{1}-z_{2})2}$. Then, using the Cauchy determinant formula (see e. g. [ABDKS4, hock2023xy]), one computes
(13) $\Omega_{n}=\prod_{i=1}^{n}\sqrt{dz_{i}^{+}dz_{i}^{-}}\;(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{z_{i}^{+}-z_{\sigma(i)}^{-}}.$
Next, assume that $x(z)=z$. Then
(14) $\mathbb{W}_{n}=\prod_{i=1}^{n}\Bigl{(}e^{u_{i}(\mathcal{S}(u_{i}\hbar\partial_{z_{i}})-1)y_{i}}dz_{i}\Bigr{)}\;(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{z_{i}+\frac{u_{i}\hbar}{2}-z_{\sigma(i)}+\frac{u_{\sigma(i)}\hbar}{2}}.$
In a similar situation with $x(z)=\log z$ one has
(15) $\mathbb{W}_{n}=\prod_{i=1}^{n}\Bigl{(}e^{u_{i}(\mathcal{S}(u_{i}\hbar z_{i}\partial_{z_{i}})-1)y_{i}}dz_{i}\Bigr{)}\;(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{e^{\frac{u_{i}\hbar}{2}}z_{i}-e^{-\frac{u_{\sigma(i)}\hbar}{2}}z_{\sigma(i)}}.$
Remark 3.8.
In this paper our sign conventions are different comparing to the previous papers including [ABDKS3, ABDKS4]. Moreover, the definition of the extended $n$-point differentials $\Omega_{n}$ is slightly different comparing to [ABDKS3], were by $\Omega_{n}$ we denote corresponding functions.
3.3. $x-y$ duality and its basic properties
Starting from a given system of differentials $\{\omega^{(g)}_{n}\}$, the $x-y$ duality construction produces a new one denoted by $\{\omega^{\vee,(g)}_{n}\}$ with the role of $x$ and $y$ functions swapped. That is, we have
(16) $(x^{\vee},y^{\vee})$ $=(y,x),$
(17) $\omega^{\vee,(0)}_{1}$ $=y^{\vee}\,dx^{\vee}=x\,dy.$
Definition 3.9.
The $x-y$ dual differentials are defined by
(18) $\omega^{\vee,(g)}_{n}(z_{\llbracket{n}\rrbracket})=(-1)^{n}[\hbar^{2g-2+n}]\prod_{i=1}^{n}\biggl{(}\sum_{r=0}^{\infty}\bigl{(}-d_{i}\tfrac{1}{dy_{i}}\bigr{)}^{r}[u_{i}^{r}]\biggr{)}\;\mathbb{W}_{n}(z_{\llbracket{n}\rrbracket},u_{\llbracket{n}\rrbracket})$
for $(g,n)\neq(0,1)$ and $\omega^{\vee,(0)}_{1}=x\,dy$, where $\mathbb{W}_{n}$ is the $n$-point differential of the previous section and $d\frac{1}{dx}$ is the operator acting in the space of meromorphic differentials and sending a given $1$-form $\eta$ to $d\frac{\eta}{dx}$.
The polynomiality property of Remark 3.5 implies that the expression for $\omega^{\vee,(g)}_{n}$ involves a finite number of summands only. For example, for small $(g,n)$ we have
(19) $\omega^{\vee,(0)}_{2}$ $=\omega^{(0)}_{2},$
(20) $\omega^{\vee,(0)}_{3}$ $=-\omega^{(0)}_{3}+d_{1}\tfrac{\omega^{(0)}_{2}(z_{1},z_{2})\omega^{(0)}_{2}(z_{1},z_{3})}{dy_{1}dx_{1}}+d_{2}\tfrac{\omega^{(0)}_{2}(z_{2},z_{3})\omega^{(0)}_{2}(z_{2},z_{1})}{dy_{2}dx_{2}}+d_{3}\tfrac{\omega^{(0)}_{2}(z_{3},z_{1})\omega^{(0)}_{2}(z_{3},z_{2})}{dy_{3}dx_{3}},$
(21) $\omega^{(0),{\rm reg}}_{2}$ $=\omega^{(0)}_{2}-\tfrac{dx_{1}dx_{2}}{(x_{1}-x_{2})^{2}},$
(22) $\omega^{\vee,(1)}_{1}$ $=-\omega^{(1)}_{1}+d\biggl{(}\frac{\omega^{(0),{\rm reg}}_{2}(z,z)}{2\,dx\,dy}-\frac{1}{24}\partial_{y}^{2}\frac{dy}{dx}\biggr{)}.$
Here are the basic properties of $x-y$ duality, for more details see [ABDKS1, ABDKS3, ABDKS4].
Theorem 3.10.
  1. 1.
    The dual differentials $\omega^{\vee,(g)}_{n}$ are meromorphic, symmetric, and possess the property that $\omega^{\vee,(g)}_{n}-\delta_{(g,n),(0,2)}\frac{dx^{\vee}_{1}dx^{\vee}_{2}}{(x^{\vee}_{1}-x^{\vee}_{2})^{2}}$ are regular on the diagonals.
  2. 2.
    This correspondence is a duality indeed: the original differentials $\omega^{(g)}_{n}$ can be recovered from the dual ones by a similar formula with the role of $x$ and $y$ functions swapped
    (23) $\omega^{(g)}_{n}(z_{\llbracket{n}\rrbracket})=(-1)^{n}[\hbar^{2g-2+n}]\prod_{i=1}^{n}\biggl{(}\sum_{r=0}^{\infty}\bigl{(}-d_{i}\tfrac{1}{dx_{i}}\bigr{)}^{r}[v_{i}^{r}]\biggr{)}\;\mathbb{W}^{\vee}_{n}(z_{\llbracket{n}\rrbracket},v_{\llbracket{n}\rrbracket}),$
    $(g,n)\not=(0,1)$, where the differentials $\mathbb{W}^{\vee}_{n}$ are obtained from the differentials $\omega^{\vee,(\tilde{g})}_{\tilde{n}}$ through the differentials $\Omega^{\vee}_{n}$ in a way similar to (8)–(9) by summation over graphs.
  3. 3.
    Moreover, the extended dual differentials $\mathbb{W}_{n}$ and $\mathbb{W}^{\vee}_{n}$ can be obtained from one another directly by the following explicit relations that avoid combinatorics of summation over graphs
    (24) $\mathbb{W}^{\vee}_{n}(z_{\llbracket{n}\rrbracket},v_{\llbracket{n}\rrbracket})$ $=(-1)^{n}\prod_{i=1}^{n}\biggl{(}e^{-v_{i}x_{i}}\sum_{r=0}^{\infty}\bigl{(}-d_{i}\tfrac{1}{dy_{i}}\bigr{)}^{r}e^{v_{i}x_{i}}[u_{i}^{r}]\biggr{)}\;\mathbb{W}_{n}(z_{\llbracket{n}\rrbracket},u_{\llbracket{n}\rrbracket})+\delta_{n,1}\frac{dy_{1}}{v_{1}\hbar},$
    (25) $\mathbb{W}_{n}(z_{\llbracket{n}\rrbracket},u_{\llbracket{n}\rrbracket})$ $=(-1)^{n}\prod_{i=1}^{n}\biggl{(}e^{-u_{i}y_{i}}\sum_{r=0}^{\infty}\bigl{(}-d_{i}\tfrac{1}{dx_{i}}\bigr{)}^{r}e^{u_{i}y_{i}}[v_{i}^{r}]\biggr{)}\;\mathbb{W}^{\vee}_{n}(z_{\llbracket{n}\rrbracket},v_{\llbracket{n}\rrbracket})+\delta_{n,1}\frac{dx_{1}}{u_{1}\hbar}.$
  4. 4.
    The (half-)differentials $\Omega_{n}$ and $\Omega^{\vee}_{n}$ are related to each other by integral transforms, defined by an asymptotic expansion of Gaussian integrals ([ABDKS3, Theorem 4.3]).
  5. 5.
    The $x-y$ duality preserves KP integrability ([ABDKS3, Theorem 2.3]). Namely, the system of differentials $\{\omega^{(g)}_{n}\}$ possesses the KP integrability property if and only if the differentials $\{\omega^{\vee,(g)}_{n}\}$ possesses the KP integrability property. More explicitly, denote
    (26) $\mathbb{K}(z^{+},z^{-})=\Omega_{1}(z^{+},z^{-}),\qquad\mathbb{K}^{\vee}(z^{+},z^{-})=\Omega^{\vee}_{1}(z^{+},z^{-}).$
    Then the following determinantal relations for the differentials $\omega^{(g)}_{n}$ hold true
    (27) $\sum_{g=0}^{\infty}\hbar^{2g-2+n}\omega^{(g)}_{n}$ $=(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\mathbb{K}(z_{i},z_{\sigma(i)}),\qquad n\geq 2,$
    (28) $\sum_{g=1}^{\infty}\hbar^{2g-1}\omega^{(g)}_{1}$ $=\lim_{z_{2}\to z_{1}}\Bigl{(}\mathbb{K}(z_{1},z_{2})-\tfrac{\sqrt{dx_{1}\,dx_{2}}}{x_{1}-x_{2}}\Bigr{)}$
    if and only if the dual differentials satisfy similar determinantal relations
    (29) $\sum_{g=0}^{\infty}\hbar^{2g-2+n}\omega^{\vee,(g)}_{n}$ $=(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\mathbb{K}^{\vee}(z_{i},z_{\sigma(i)}),\qquad n\geq 2,$
    (30) $\sum_{g=1}^{\infty}\hbar^{2g-1}\omega^{\vee,(g)}_{1}$ $=\lim_{z_{2}\to z_{1}}\Bigl{(}\mathbb{K}^{\vee}(z_{1},z_{2})-\tfrac{\sqrt{dy_{1}\,dy_{2}}}{y_{1}-y_{2}}\Bigr{)}.$
    Moreover, in the KP integrable case the differentials $\Omega_{n}$ and their $x-y$ dual differentials $\Omega^{\vee}_{n}$ admit determinantal presentations
    (31) $\Omega_{n}(z^{+}_{\llbracket n\rrbracket},z^{-}_{\llbracket n\rrbracket})$ $=(-1)^{n-1}\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\Omega_{1}(z^{+}_{i},z^{-}_{\sigma(i)}),$ $n\geq 1,$
    (32) $\Omega^{\vee}_{n}(z^{+}_{\llbracket n\rrbracket},z^{-}_{\llbracket n\rrbracket})$ $=(-1)^{n-1}\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\Omega^{\vee}_{1}(z^{+}_{i},z^{-}_{\sigma(i)}),$ $n\geq 1.$
  6. 6.
    Let $p$ be a zero of $dx$ which is not a zero of $dy$. Then all differentials $\omega^{\vee,(g)}_{n}$, $2g-2+n\geq 0$, are regular at $p$ if and only if the differentials $\omega^{(g)}_{n}$ satisfy the loop equations at $p$.
    Vice versa, let $p$ be a zero of $dy$ which is not a zero of $dx$. Then all differentials $\omega^{(g)}_{n}$, $2g-2+n\geq 0$, are regular at $p$ if and only if the differentials $\omega^{\vee,(g)}_{n}$ satisfy the loop equations at $p$.
  7. 7.
    The $x-y$ duality preserves (logarithmic) topological recursion. Namely, let $dx$ and $dy$ have non-intersecting sets of zeros, all of which are simple, and $\omega^{(0)}_{2}=\omega^{\vee,(0)}_{2}$ be the Bergman kernel, that is, it has no singularities away from the diagonal and its ${\mathfrak{A}}$-periods are vanishing for some choice of collection of $({\mathfrak{A}},{\mathfrak{B}})$ cycles on $\Sigma$. If the functions $x$ and $y$ themselves are meromorphic, then the differentials $\omega^{(g)}_{n}$ solve topological recursion for the spectral curve data $(x,y)$ if and only if the differentials $\omega^{\vee,(g)}_{n}$ solve topological recursion for the spectral curve data $(y,x)$.
    More generally, if the differentials $dx$ and $dy$ are meromorphic, then the differentials $\omega^{(g)}_{n}$ solve logarithmic topological recursion for the spectral curve data $(x,y)$ if and only if the differentials $\omega^{\vee,(g)}_{n}$ solve logarithmic topological recursion for the spectral curve data $(y,x)$.
Example 3.11.
We demonstrate here one relatively simple but quite shining example of $x-y$ duality. More examples and applications can be found in [ABDKS1, ABDKS3, ABDKS4]. Consider a genus zero spectral curve with the spectral curve data
(33) $x=z,\qquad y=\log z.$
Both $dx=dz$ and $dx^{\vee}=dy=\frac{dz}{z}$ have no zeros and both topological recursions for this spectral curve and its dual are trivial. However, in the presence of logarithmic singularities, it is more natural to consider the logarithmic topological recursion. The LogTR differentials for this spectral curve and its dual are given by
(34) $\omega^{(g)}_{1}$ $=\Bigl{(}[\hbar^{2g}]\tfrac{1}{\mathcal{S}(\hbar\partial_{z})}\log z\Bigr{)}\;dz,$
(35) $\omega^{\vee,(0)}_{1}$ $=z\;d\log z,$
(36) $\omega^{(0)}_{2}$ $=\omega^{\vee,(0)}_{2}=\frac{dz_{1}dz_{2}}{(z_{1}-z_{2})^{2}},$
and all other differentials $\omega^{(g)}_{n}$ and $\omega^{\vee,(g)}_{n}$ are equal to zero (here we omit the pre-superscript log in the notation). Respectively, analogously to (14)–(15) we obtain explicitly
(37) $\mathbb{W}_{n}$ $=\prod_{i=1}^{n}\Bigl{(}e^{u_{i}\bigl{(}\frac{\mathcal{S}(u_{i}\hbar\partial_{z_{i}})}{\mathcal{S}(\hbar\partial_{z_{i}})}-1\bigr{)}\log z_{i}}dz_{i}\Bigr{)}\;(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{z_{i}+\frac{u_{i}\hbar}{2}-z_{\sigma(i)}+\frac{u_{\sigma(i)}\hbar}{2}},$
(38) $\mathbb{W}^{\vee}_{n}$ $=\prod_{i=1}^{n}\Bigl{(}e^{v_{i}(\mathcal{S}(v_{i}\hbar)-1)z_{i}}dz_{i}\Bigr{)}\;(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{e^{\frac{v_{i}\hbar}{2}}z_{i}-e^{-\frac{v_{\sigma(i)}\hbar}{2}}z_{\sigma(i)}}.$
These differentials are related to one another by (24)–(25), which are nontrivial identities even for this elementary case.
4. Symplectic duality
4.1. Definition of symplectic duality
Consider a spectral curve data $(\Sigma,x,y)$ and a function $\psi(\theta)$ defined on the rational curve. Define a new spectral curve data $(x^{\dagger},y^{\dagger})$ by
(39) $x^{\dagger}(z)=x(z)+\psi(y(z)),\qquad y^{\dagger}(z)=y(z),$
where $z$ is a point on $\Sigma$. This correspondence is a duality in a sense that the inverse transformation is produced by the dual function $\psi^{\dagger}(\theta)=-\psi(\theta)$,
(40) $x(z)=x^{\dagger}(z)+\psi^{\dagger}(y(z)).$
The symplectic duality construction presented below involves not only the function $\psi$ but also some its $\hbar$-deformation $\tilde{\psi}(\theta,\hbar)$ which is a power series in 2superscriptPlanck-constant-over-$\hbar^{2}$ such that $\tilde{\psi}(\theta,0)=\psi(\theta)$. The analytic properties of $\psi$ and $\tilde{\psi}$ will be discussed later. Basically, for the following definition to make sense we require that $\psi^{\prime}(y(z))$ and $[\hbar^{2g}]\tilde{\psi}(y(z),\hbar)$ for $g\geq 1$ are meromorphic.
Definition 4.1.
Given a system of symmetric differentials $\{\omega^{(g)}_{n}\}$ such that $\omega^{(0)}_{1}=y\,dx$, the symplectic duality correspondence produces a new system of symmetric differentials denoted by $\{\omega^{{\dagger},(g)}_{n}\}$ and given by the following relations:
(41) $\omega^{{\dagger},(0)}_{1}$ $=y\,dx^{\dagger},$
(42) $\omega^{{\dagger},(g)}_{n}(z_{\llbracket{n}\rrbracket})$ $=[\hbar^{2g-2+n}]U_{1}\dots U_{n}\mathbb{W}_{n}(z_{\llbracket{n}\rrbracket},u_{\llbracket{n}\rrbracket}),\quad n\geq 2,$
where $\mathbb{W}_{n}$ is the extended differential of Definition 3.2 and $U_{i}$ is the operator acting on differentials in $z=z_{i}$ depending on an additional variable $u=u_{i}$ in a polynomial way defined by
(43) $UF$ $=\sum_{j,r\geq 0}\bigl{(}-d\tfrac{1}{dx^{\dagger}}\bigr{)}^{j}[v^{j}]\Bigl{(}\mathop{\big{\lfloor}_{{\theta}\to{y}}}e^{-v\psi(\theta)}\partial_{\theta}^{r}e^{v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)}\Bigr{)}[u^{r}]F(z,u)$
$=\sum_{j,r\geq 0}\bigl{(}d\tfrac{1}{dx^{\dagger}}\bigr{)}^{j}[v^{j}]\Bigl{(}\mathop{\big{\lfloor}_{{\theta}\to{y}}}e^{v\psi(\theta)}\partial_{\theta}^{r}e^{-v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)}\Bigr{)}[u^{r}]F(z,u).$
In the case $n=1$ the formula is a bit more involved:
(44) $\omega^{{\dagger},(g)}_{1}(z_{1})$ $=[\hbar^{2g-1}]U_{1}\mathbb{W}_{1}(z_{1},u_{1})+[\hbar^{2g}]\sum_{j=1}^{\infty}\bigl{(}d_{1}\tfrac{1}{dx^{\dagger}_{1}}\bigr{)}^{j-1}[v^{j}]\left(e^{-v(\mathcal{S}(v\hbar\partial_{y_{1}})\tilde{\psi}(y_{1})-\psi(y_{1}))}dy_{1}\right).$
4.2. Example: hypergeometric tau functions and associated differentials
Let $\psi(\theta)$ and $y(z)$ be some formal power series such that $\psi(0)=0$, $y(0)=0$. Assume that we are given also some 2superscriptPlanck-constant-over-$\hbar^{2}$-deformations $\hat{\psi}(\theta,\hbar)$ and $\hat{y}(z,\hbar)$ of these functions. These data determine the so-called Orlov–Scherbin (or, in other terminology, hypergeometric) tau functions [OS] and the corresponding $n$-point differentials [BDKS1, BDKS2].
Definition 4.2.
The Orlov–Scherbin (or hypergeometric) tau function is given by the sum over partitions
(45) $Z^{\mathrm{OS}}=\sum_{\lambda}{}_{\lambda}(t){}_{\lambda}(\hbar^{-1}s)\exp\left(\sum_{(i,j)\in\lambda}\hat{\psi}(\hbar(i-j),\hbar)\right),$
where (${}_{\lambda}(t)$ is the Schur function in the variables $t=(t_{1},t_{2},\dots)$ labeled by the partition $\lambda$ and the parameters $s_{i}=s_{i}(\hbar)$ for the second Schur function are the expansion coefficients defined by $\hat{y}(z,\hbar)=\sum_{j=1}^{\infty}s_{j}(\hbar)z^{j}$.
Definition 4.3.
The Orlov–Scherbin differentials are defined as
(46) $\sum_{g=0}^{\infty}\hbar^{2g-2+n}\omega^{\mathrm{OS},(g)}_{n}=\prod_{j=1}^{\infty}\mathop{\big{\lfloor}_{{t_{j}}\to{0}}}\prod_{i=1}^{n}\bigg{(}d_{i}\sum_{k_{i}=1}^{\infty}X_{i}^{k_{i}}\partial_{t_{k_{i}}}\bigg{)}\log Z^{\mathrm{OS}}.$
Under some natural analyticity assumptions (see [BDKS2, Definition 1.4]) on the initial data, the Orlov–Scherbin differentials extend as global meromorphic $n$-differentials on the spectral curve $\Sigma=\mathbb{C}P^{1}$ and we have
(47) $X$ $=z\,e^{-\psi(y(z))},$
(48) $\omega^{{\rm OS},(0)}_{1}$ $=y\frac{dX}{X}.$
Let us denote
(49) $x=\log X=\log z-\psi(y(z)),\quad x^{\dagger}=\log z,\quad\tilde{\psi}(\theta)=\frac{1}{\mathcal{S}(\hbar\partial_{\theta})}\hat{\psi}(\theta).$
Then we have $\omega^{{\rm OS},(0)}_{1}=y\,dx$ and the following proposition is a reformulation of [BDKS2, Proposition 2.3].
Proposition 4.4.
The Orlov–Scherbin differentials $\omega^{(g)}_{n}=\omega^{{\rm OS},(g)}_{n}$ can be computed as symplectic dual to the trivial system of differentials
(50) $\omega^{{\dagger},(g)}_{n}=\delta_{n,1}[\hbar^{2g}]\hat{y}_{1}dx_{1}+\delta_{(g,n),(0,2)}\tfrac{dz_{1}dz_{2}}{(z_{1}-z_{2})^{2}}$
associated with the spectral curve $(\log z,y(z))$ and computed with respect to the function $\tilde{\psi}^{\dagger}(\theta,\hbar)=-\tilde{\psi}(\theta,\hbar)=-\frac{1}{\mathcal{S}(\hbar\partial_{\theta})}\hat{\psi}(\theta,\hbar)$.
More explicitly, the formula for the differentials of Orlov–Scherbin family for $g\geq 0$, $n\geq 1$ reads
(51) $\omega^{(g)}_{n}$ $=[\hbar^{2g-2+n}]\prod_{i=1}^{n}\biggl{(}\sum_{j,r\geq 0}\bigl{(}d_{i}\tfrac{1}{dx_{i}}\bigr{)}^{j}\Bigl{(}\mathop{\big{\lfloor}_{{\theta}\to{y_{i}}}}[v^{j}]e^{-v\psi(\theta)}\partial_{\theta}^{r}e^{v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)}\Bigr{)}[u_{i}^{r}]\biggr{)}$
$\qquad\prod_{i=1}^{n}\Bigl{(}e^{u_{i}(\mathcal{S}(u_{i}\hbar\,z_{i}\partial_{z_{i}})\hat{y}_{i}-y_{i})}dz_{i}\Bigr{)}\cdot(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{e^{\frac{u_{i}\hbar}{2}}z_{i}-e^{-\frac{u_{\sigma(i)}\hbar}{2}}z_{\sigma(i)}}$
$\quad+\delta_{n,1}[\hbar^{2g}]\sum_{j=0}^{\infty}\bigl{(}d_{1}\tfrac{1}{dx_{1}}\bigr{)}^{j-1}[v^{j}]\left(e^{v(\mathcal{S}(v\hbar\partial_{y_{1}})\tilde{\psi}(y_{1})-\psi(y_{1}))}dy_{1}\right).$
Proof.
We refer the reader to [BDKS2, Proposition 2.3]. The only difference between the expression given there and Equation (51) is that the former uses the following expression instead of the second factor in the second line of (51):
(52) $\prod_{i=1}^{n}\frac{1}{u_{i}\hbar\mathcal{S}(u_{i}\hbar)z_{i}}\sum_{\gamma\in\Gamma_{n}}\prod_{(v_{k},v_{l})\in E_{\gamma}}w_{k,l},$
where the sum is taken over all simple connected graphs $\gamma$ on $n$ labeled vertices, and the edge $(v_{k},v_{l})$ in the set of edges $E_{\gamma}$ of $\gamma$ is decorated by
(53) $w_{k,\ell}$ $\coloneqq e^{\hbar^{2}u_{k}u_{l}\mathcal{S}(u_{k}\hbar Q_{k}D_{k})\mathcal{S}(u_{\ell}\hbar Q_{\ell}D_{\ell})\frac{z_{k}z_{\ell}}{(z_{k}-z_{\ell})^{2}}}-1.$
The identification of this expression with
(54) $(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{e^{\frac{u_{i}\hbar}{2}}z_{i}-e^{-\frac{u_{\sigma(i)}\hbar}{2}}z_{\sigma(i)}}$
is the connected version of the Cauchy determinant formula
(55) $(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{a_{i}-b_{\sigma(i)}}=\prod_{i=1}^{n}\frac{1}{a_{i}-b_{i}}\;\sum_{\gamma\in\Gamma_{n}}\prod_{(v_{k},v_{\ell})\in E_{\gamma}}\Bigl{(}\tfrac{(a_{k}-a_{\ell})(b_{k}-b_{\ell})}{(a_{k}-b_{\ell})(b_{k}-a_{\ell})}-1\Bigr{)}$
applied to $a_{k}=e^{\frac{u_{k}\hbar}{2}}z_{k}=e^{\frac{u_{k}\hbar}{2}z_{k}\partial_{z_{k}}}z_{k}$, $b_{k}=e^{-\frac{u_{k}\hbar}{2}}z_{k}=e^{-\frac{u_{k}\hbar}{2}z_{k}\partial_{z_{k}}}z_{k}$.
In the case $n=1$, $g\geq 1$ the expression for $\omega^{(g)}_{n}$ contains an additional term, see [BDKS2, Equation (32)], which matches with the last summand in (51). ∎
4.3. Example: $(t,s)$-weighted hypermaps
To a given series $\phi(\theta)$ such that $\phi(0)=1$ one associates a tau function depending on two sets of times $t=(t_{1},t_{2},\dots)$ and $s=(s_{1},s_{2},\dots)$ and a two-index set of $(m,n)$-point functions $H^{(g)}_{m,n}(X_{1},\dots,X_{m},Y_{1},\dots,Y_{n})$ (depending also on $(t,s)$-parameters). In the case $\phi(\theta)=1+\theta$ these functions enumerate genus $g$ hypermaps with marked and unmarked vertices of two colors known also as cumulants of 2-matrix model, see [BDKS4, Introduction].
More precisely, we use the following modification of the construction given in the previous section. We consider
(56) $Z=\sum_{\lambda}{}_{\lambda}(\hbar^{-1}t){}_{\lambda}(\hbar^{-1}s)\prod_{(i,j)\in\lambda}\phi(\hbar(i-j)),$
for $\phi(\theta)\coloneqq 1+\theta$ and we define $H^{(g)}_{m,n}$ as
(57) $\sum_{g=0}^{\infty}\hbar^{2g-2}H^{(g)}_{m,n}=\prod_{i=1}^{m}\bigg{(}d_{i}\sum_{k_{i}=1}^{\infty}X_{i}^{k_{i}}\partial_{t_{k_{i}}}\bigg{)}\prod_{i=1}^{n}\bigg{(}d_{i}\sum_{\ell_{i}=1}^{\infty}Y_{i}^{\ell_{i}}\partial_{s_{\ell_{i}}}\bigg{)}\log Z.$
For a general $\phi$ they enumerate certain weighted hypermaps or weighted Hurwitz numbers, see e.g. [ALS, Harnad].
Consider the case when $\phi$ is a polynomial and in $H^{(g)}_{m,n}$ we set all but finitely many $(t,s)$ parameters to zero. Under these assumptions, one constructs in [BDKS4] a rational spectral curve $\Sigma=\mathbb{C}\mathrm{P}^{1}$ equipped with three rational functions $X,Y,\Theta$ satisfying $X\,Y\,\phi(\Theta)=1$ such that $X$ is a local coordinate at the point $z=0$, and $Y$ is a local coordinate at the point $z=\infty$. It is proved that the function $H^{(g)}_{m,n}$ introduced initially as a power series in $X,Y$ variables extends as a global rational function on $\Sigma^{m+n}$. Denote
(58) $x$ $=\log X,\qquad x^{\dagger}=-\log Y,\qquad y=\Theta,$
(59) $\psi(\theta)$ $=\log\phi(\theta),$
(60) $\omega^{(0)}_{1}$ $=\Theta\frac{dX}{X}=y\,dx,\qquad\omega^{{\dagger},(0)}_{1}=-\Theta\frac{dY}{Y}=y\,dx^{\dagger},$
(61) $\omega^{(0)}_{2}$ $=\omega^{{\dagger}\,(0)}_{2}=\frac{dz_{1}dz_{2}}{(z_{1}-z_{2})^{2}},$
(62) $\omega^{(g)}_{n}$ $=d_{1}\dots d_{n}H^{(g)}_{n,0},\quad 2g-2+n>0,$
(63) $\omega^{{\dagger},(g)}_{n}$ $=(-1)^{n}d_{1}\dots d_{n}H^{(g)}_{0,n},\quad 2g-2+n>0.$
With this notation, the results of [BDKS4] expressing $H^{(g)}_{n,0}$ via $H^{(h)}_{0,m}$ can be reformulated in the following way
Theorem 4.5.
The differentials $\omega^{(g)}_{n}$, $\omega^{{\dagger},(g)}_{n}$ are symplectic dual for the spectral curves $(\log X,\Theta)$ and $(-\log Y,\Theta)$ and computed with respect to the $\hbar$-deformation $\tilde{\psi}(\theta)=\frac{1}{\mathcal{S}(\hbar\theta)}\log\phi(\theta)$ of the function $\log\phi$.
Remark 4.6.
In fact, no explicit expression of $H^{(g)}_{n,0}$ via $H^{(h)}_{0,m}$ is given explicitly in [BDKS4]. In loc. cit. only the formulas expressing $H^{(g)}_{m,n+1}$ through $H^{(g^{\prime})}_{m^{\prime},n^{\prime}}$ with $n^{\prime}\leq n$ are given. But then one can combine these expressions into a formula expressing $H^{(g)}_{n,0}$ via $H^{(h)}_{0,m}$.
4.4. Example: $\psi=\log(\theta)$ symplectic duality as the original formulation of $x-y$ duality
Consider the initial differentials of $x-y$ duality
(64) $-\omega^{(0)}_{1}=-y\,dx=xy\;d\log x^{-1},\qquad\omega^{\vee,(0)}_{1}=x\,dy=xy\;d\log y.$
We see that these differentials can also be treated as initial differentials for the spectral curves $(\log x^{-1},x\,y)$ and $(\log y,x\,y)$, respectively. Note that $\log y-\log x^{-1}=\psi(x\,y)$ with $\psi(\theta)=\log\theta$. This motivates the following result.
Proposition 4.7.
The dual differentials $\omega^{\vee,(g)}$ of $x-y$ duality coincide with the symplectic dual of the differentials $(-1)^{n}\omega^{(g)}_{n}$ computed with respect to symplectic dual spectral curves $(\log x^{-1},x\,y)$ and $(\log y,x\,y)$ and the $\hbar$-deformation
(65) $\tilde{\psi}(\theta)=\frac{1}{\mathcal{S}(\hbar\partial_{\theta})}\log\theta$
of the function $\psi(\theta)=\log\theta$.
This proposition is proved in [ABDKS1, Section 3]. This treatment of $x-y$ duality appeared in the literature as its first known formula, see [borot2023functional, BDKS-fullysimple, ABDKS1]. Note that this definition can only be applied if both $x$ and $y$ are meromorphic. Besides, it is not invariant under a change of $x$ or $y$ by additive constants. Meanwhile, for the correctness of Definition 3.9 it is sufficient to assume that $dx$ and $dy$ are meromorphic and the duality transformation is uniquely determined by these differentials.
5. Symplectic duality as a combination of $x-y$ dualities
Consider the spectral curves $x-y$ dual to the original one given by the pair of functions $(x,y)$ and its symplectic dual one $(x^{\dagger},y^{\dagger})=(x+\psi(y),y)$,
(66) $(x^{\vee},y^{\vee})$ $=(y,x),$ $\omega^{\vee,(0)}_{1}=x\,dy,$
(67) $(x^{{\dagger}\vee},y^{{\dagger}\vee})$ $=(y,x+\psi(y)),$ $\omega^{{\dagger}\vee,(0)}_{1}=(x+\psi(y))\,dy.$
These curves are very close. For example, if $x$ and $y$ are meromorphic and $\tilde{\psi}=\psi$ is rational, then the topological recursion differentials produced by these spectral curves are the same for $(g,n)\neq(0,1)$. This provides a hint that the $x-y$ dual differentials of the original system of differentials and its symplectic dual should be related in a more general formal setting. The following theorem claims that it is indeed the case.
Theorem 5.1.
Consider a given collection of differentials $\{\omega^{(g)}_{n}\}$ and its symplectic dual one $\{\omega^{{\dagger},(g)}_{n}\}$ computed with respect to a chosen function $\tilde{\psi}$. Then the $x-y$ dual of these two systems of differentials, $\{\omega^{\vee,(g)}_{n}\}$, and $\{\omega^{{\dagger}\vee,(g)}_{n}\}$, respectively, are related to one another by
(68) $\omega^{{\dagger}\vee,(g)}_{n}=\omega^{\vee,(g)}_{n}+\delta_{n,1}\Bigl{(}[\hbar^{2g}]\tilde{\psi}(y(z),\hbar)\Bigr{)}\,dy.$
In other words, the symplectic dual differentials can be obtained by the following combination of three transformations:
  1. (a)
    apply $x-y$ duality to the original system of differentials;
  2. (b)
    modify $(g,1)$-differentials of the obtained system by the summands prescribed by (68);
  3. (c)
    apply $x-y$ duality once again.
Proof.
In fact, while applying these transformations, there is no need to involve combinatorics of summation over graphs each time. It is much more instructive to follow the action of these transformations on the corresponding extended differentials which, according to Definition 3.2, already contain these summations.
In particular, Equation (68) rewritten in terms of extended differentials reads
(69) $\mathbb{W}^{{\dagger}\vee}_{n}(z_{\llbracket{n}\rrbracket},v_{\llbracket{n}\rrbracket})=\mathbb{W}^{\vee}_{n}(z_{\llbracket{n}\rrbracket},v_{\llbracket{n}\rrbracket})\prod_{i=1}^{n}e^{v_{i}(\mathcal{S}(v_{i}\hbar\partial_{y_{i}})\tilde{\psi}(y_{i})-\psi(y_{i}))}.$
Substituting to (23)–(24) we obtain the following expression for the dual differentials for $n\geq 2$ suggested by Theorem 5.1
(70) $\omega^{{\dagger},(g)}_{n}=[\hbar^{2g-2+n}]\prod_{i=1}^{n}\Bigl{(}\sum_{r,j\geq 0}\bigl{(}-d_{i}\tfrac{1}{dx^{\dagger}_{i}}\bigr{)}^{j}[v_{i}^{j}]e^{v_{i}(\mathcal{S}(v_{i}\hbar\partial_{y_{i}})\tilde{\psi}(y_{i})-\psi(y_{i}))}e^{-v_{i}x_{i}}\bigl{(}-d_{i}\tfrac{1}{dy_{i}}\bigr{)}^{r}e^{v_{i}x_{i}}[u_{i}^{r}]\Bigr{)}\mathbb{W}_{n},$
and expressions for $\omega^{{\dagger},(g)}_{1}$ differentials contain correction which are easy to trace.
Comparing with the actual definition (42) of the differentials $\omega^{{\dagger},(g)}_{n}$ (and its version for the special case $n=1$ given in (44)) we see that Theorem follows from the following operator identity: for any meromorphic differential $F(z,u)$ depending on $u$ in polynomial way we have
(71) $UF$ $=\sum_{r,j\geq 0}\bigl{(}-d\tfrac{1}{dx^{\dagger}}\bigr{)}^{j}[v^{j}]\Bigl{(}\mathop{\big{\lfloor}_{{\theta}\to{y}}}e^{-v\psi(\theta)}\partial_{\theta}^{r}e^{v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)}\Bigr{)}[u^{r}]F(z,u)$
$=\sum_{r,j\geq 0}\bigl{(}-d\tfrac{1}{dx^{\dagger}}\bigr{)}^{j}[v^{j}]e^{v(\mathcal{S}(v\hbar\partial_{y})\tilde{\psi}(y)-\psi(y))}e^{-vx}\bigl{(}-d\tfrac{1}{dy}\bigr{)}^{r}e^{vx}[u^{r}]F(z,u).$
We claim that not only the whole expressions on both sides but also their summands for each particular $r$ agree. Namely, denote
(72) $A(z,v)=e^{-v\psi(y)}[u^{r}]\frac{F(z,u)}{dy},\qquad B(z,v)=e^{v\mathcal{S}(v\hbar\partial_{y})\tilde{\psi}(y)}.$
With this notation and using equalities $x+\psi(y)=x^{\dagger}$, $\bigl{(}d\tfrac{1}{dx}\bigr{)}^{j}=dx\,\partial_{x}^{j}\frac{1}{dx}$, we can write the equality of the $r$th terms in (71) that we are going to prove in the following form
(73) $\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}\;\frac{dy}{dx^{\dagger}}[v^{j}]A\partial_{y}^{r}B=\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}\;\frac{dy}{dx^{\dagger}}[v^{j}]Be^{-vx^{\dagger}}(-\partial_{y})^{r}e^{vx^{\dagger}}A.$
Lemma 5.2.
Identity (73) holds for any $r\geq 0$ and for any functions $x^{\dagger}(z)$, $y(z)$, $A(z,v)$, $B(z,v)$, where $A$ and $B$ considered as functions in $v$ are power series such that the products $A\,B$ and $A\partial_{z}B$ are polynomial in $v$.
Proof.
Indeed, using that
(74) $0=\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}[v^{j}](\partial_{x^{\dagger}}+v)A\,B=\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}[v^{j}]e^{-vx^{\dagger}}\,\partial_{x^{\dagger}}\,e^{vx^{\dagger}}A\,B\\ =\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}\frac{dy}{dx^{{\dagger}}}[v^{j}]e^{-vx^{\dagger}}\,\partial_{y}\,e^{vx^{\dagger}}A\,B=\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}\frac{dy}{dx^{{\dagger}}}[v^{j}](A\,\partial_{y}B+e^{-vx^{{\dagger}}}B\,\partial_{y}(e^{vx^{{\dagger}}}A))$
we conclude that
(75) $\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}\frac{dy}{dx^{{\dagger}}}[v^{j}]A\,\partial_{y}B=\sum_{j=0}^{\infty}(-\partial_{x^{\dagger}})^{j}\frac{dy}{dx^{{\dagger}}}[v^{j}]e^{-vx^{{\dagger}}}B\,(-\partial_{y})(e^{vx^{{\dagger}}}A).$
Applying this relation $r$ times we obtain the equality of Lemma. ∎
This Lemma proves (71) and hence Theorem 5.1. ∎
Remark 5.3.
By polynomiality here we mean that each coefficient of the $\hbar$ expansion of the expression is a polynomial.
As an immediate corollary of Theorem 5.1 we conclude that the symplectic duality transformation shares most of the properties valid for $x-y$ duality, see Theorem 3.10.
Theorem 5.4.
1. The symplectic dual differentials $\omega^{{\dagger},(g)}_{n}$ are meromorphic, symmetric, and possess the property that $\omega^{{\dagger},(g)}_{n}-\delta_{(g,n),(0,2)}\frac{dx^{\dagger}_{1}dx^{\dagger}_{2}}{(x^{\dagger}_{1}-x^{\dagger}_{2})^{2}}$ are regular on the diagonals.
2. This correspondence is a duality indeed: the original differentials $\omega^{(g)}_{n}$ can be recovered from the symplectic dual ones by a dual formula that in the case $n\geq 2$ reads
(76) $\omega^{(g)}_{n}(z_{\llbracket{n}\rrbracket})=$ $[\hbar^{2g-2+n}]U^{\dagger}_{1}\dots U^{\dagger}_{n}\mathbb{W}^{\dagger}_{n}(z_{\llbracket{n}\rrbracket},w_{\llbracket{n}\rrbracket}),$
where
(77) $U^{\dagger}F$ $=\sum_{j,r\geq 0}\bigl{(}-d\tfrac{1}{dx}\bigr{)}^{j}[v^{j}]\Bigl{(}\mathop{\big{\lfloor}_{{\theta}\to{y}}}e^{v\psi(\theta)}\partial_{\theta}^{r}e^{-v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)}\Bigr{)}[w^{r}]F(z,w),$
and for $n=1$ we have
(78) $\omega^{{\dagger},(g)}_{1}(z_{1})$ $=[\hbar^{2g-1}]U^{\dagger}_{1}\mathbb{W}^{\dagger}_{1}(z_{1},u_{1})+[\hbar^{2g}]\sum_{j=1}^{\infty}\bigl{(}d_{1}\tfrac{1}{dx_{1}}\bigr{)}^{j-1}[v^{j}]\left(e^{v(\mathcal{S}(v\hbar\partial_{y_{1}})\tilde{\psi}(y_{1})-\psi(y_{1}))}d_{1}y_{1}\right).$
3. Moreover, the extended dual differentials $\mathbb{W}_{n}$ and $\mathbb{W}^{\dagger}_{n}$ can be obtained from one another directly by the following explicit relations that avoid combinatorics of summation over graphs
(79) $\mathbb{W}^{{\dagger}}_{n}(z_{\llbracket{n}\rrbracket},w_{\llbracket{n}\rrbracket})$ $=\tilde{U}_{1}\dots\tilde{U}_{n}\mathbb{W}(z_{\llbracket{n}\rrbracket},u_{\llbracket{n}\rrbracket})+\delta_{n,1}\frac{dx^{\dagger}_{1}}{\hbar w_{1}},$
(80) $\tilde{U}F$ $=e^{-wy}\sum_{j,r\geq 0}\bigl{(}-d\tfrac{1}{dx^{\dagger}}\bigr{)}^{j}[v^{j}]\Bigl{(}\mathop{\big{\lfloor}_{{\theta}\to{y}}}e^{-v\psi(\theta)}\partial_{\theta}^{r}e^{v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)+w\theta}\Bigr{)}[u^{r}]F(z,u),$
(81) $\mathbb{W}_{n}(z_{\llbracket{n}\rrbracket},u_{\llbracket{n}\rrbracket})$ $=\tilde{U}_{1}^{\dagger}\dots\tilde{U}_{n}^{\dagger}\mathbb{W}(z_{\llbracket{n}\rrbracket},w_{\llbracket{n}\rrbracket})+\delta_{n,1}\frac{dx_{1}}{\hbar u_{1}},$
(82) $\tilde{U}^{\dagger}F$ $=e^{-uy}\sum_{j,r\geq 0}\bigl{(}-d\tfrac{1}{dx}\bigr{)}^{j}[v^{j}]\Bigl{(}\mathop{\big{\lfloor}_{{\theta}\to{y}}}e^{v\psi(\theta)}\partial_{\theta}^{r}e^{-v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)+uy}\Bigr{)}[w^{r}]F(z,w).$
The differentials $\Omega_{n}$ and $\Omega^{\dagger}_{n}$ are related to each other by integral transforms, defined by an asymptotic expansion around Gaussian integrals. For each of $2n$ variables a transform is given by a double integral (instead of single one in the $x-y$ swap case).
4. The symplectic duality preserves KP integrability. Namely, the system of differentials $\{\omega^{(g)}_{n}\}$ possesses the KP integrability property if and only if the differentials $\{\omega^{{\dagger},(g)}_{n}\}$ possesses the KP integrability property. Equivalently, the determinantal expressions for the differentials hold on one side of symplectic duality if and only if they hold on the other side of duality.
5. Assume that $\psi$ is unramified and regular at the critical values of $y$. Let $p$ be a zero of $dx$ which is not a zero of $dy$. Then all differentials $\omega^{{\dagger},(g)}_{n}$, $(g,n)\neq(0,1)$, are regular at $p$ if and only if the differentials $\omega^{(g)}_{n}$ satisfy the loop equations at $p$.
Vice versa, let $p$ be a zero of $dy$ which is not a zero of $dx$. Then all differentials $\omega^{(g)}_{n}$, $(g,n)\neq(0,1)$, are regular at $p$ if and only if the differentials $\omega^{{\dagger},(g)}_{n}$ satisfy loop equations at $p$.
All these properties are direct corollaries of Theorem 5.1 and the corresponding properties of $x-y$ duality. Part of these properties was already observed though stated a bit differently in [ABDKS2]. A slightly more delicate question on the compatibility with the projection property and its logarithmic version, and hence with topological recursion or logarithmic topological recursion will be discussed in the next section.
We finish this section with the following observation.
Proposition 5.5.
The symplectic duality possesses the following group property: the composition of two symplectic duality transformations computed with respect to the given functions $\tilde{\psi}_{1}$ and $\tilde{\psi}_{2}$, respectively, is equivalent to symplectic duality transformation computed with respect to the function $\tilde{\psi}_{1}+\tilde{\psi}_{2}$.
Proof.
Indeed, after conjugation with $x-y$ duality the transformation of symplectic duality results in a change of the $1$-point differentials by a summand proportional to $\tilde{\psi}(y)$. Such transformation obviously satisfies the group property. ∎
6. Compatibility with topological recursion
6.1. Topological recursion
The following statement is a direct corollary of Theorem 5.1.
Theorem 6.1.
Assume that $x$ and $y$ are meromorphic and $\psi$ is rational and regular at the critical values of $y$. Let $x^{\dagger}=x+\psi(y)$ and assume that all zeros of $dx$, $dy$, and $dx^{\dagger}$ are simple, and that the zeros of $dx$ are disjoint from the zeros of $dy$.
Let $\tilde{\psi}=\psi$. Then the system of differentials $\{\omega^{(g)}_{n}\}$ solves TR for the spectral curve $(\Sigma,x,y)$ if and only if its symplectic dual system of differentials $\{\omega^{{\dagger},(g)}_{n}\}$ solves TR for the spectral curve $(\Sigma,x^{\dagger},y)$.
Proof.
The notation in this proof follows the notation of Theorem 5.1 (in particular, see (66) and (67)). Note that transformation (a) of Theorem 5.1 preserves topological recursion by Theorem 3.10 (part 7.).
Let us see that after the application of transformation (b) of Theorem 5.1 TR still holds (this is a very well-known fact in the theory of topological recursion, see [EO]; still we discuss its proof here as it will be useful for our discussion of the case of LogTR below).
Keeping in mind that $\tilde{\psi}=\psi$, formula (68) says that $\omega^{\dagger,\vee,(g)}_{n}$ differs from $\omega^{\vee,(g)}_{n}$ only in $(g,n)=(0,1)$ term. The linear loop equations still hold due the fact that $\psi(\theta)$ is regular at the critical values of $x^{\vee}$ and thus $\psi(x^{\vee}(z))dx^{\vee}$ is regular and vanishing at the critical points of $x^{\vee}$. Regarding the quadratic loop equations, let us use the form from Remark 2.3, as the linear loop equations are already established. As the change in (68) affects only the $n=1$ terms then only the second summand in the quadratic loop equation (3) is affected. Note that we then have either $\Delta_{i,w}$ or $\Delta_{i,w}$ acting on $\psi(x^{\vee}(z))dx^{\vee}$, but $x^{\vee}$ is by definition symmetric under the deck transformations and thus $\Delta$ kills $\psi(x^{\vee}(z))dx^{\vee}$. This means that the quadratic loop equations are preserved. The projection property is imposed only upon the $02g-2+n>$ 0 differentials and thus it is preserved as well, as only $\omega^{(0)}_{1}$ gets changed.
Thus, after transformation (b) TR still holds, and after transformation (c) it holds as well again due to Theorem 3.10 (part 7.). In the latter statement we have used the fact that the zeros of $dx^{\dagger}$ are disjoint from the zeros of $dy$ since at the zeros of $dy$ we have $dx=dx^{\dagger}$ due to the condition that $\psi$ is regular at the critical values of $y$. ∎
Remark 6.2.
Let us stress that the above theorem works for an arbitrary compact Riemann surface $\Sigma$ of any genus.
6.2. Logarithmic topological recursion
We would like to extend the result of Theorem 6.1 to the situation when $x$ and $y$ may possess logarithmic singularities, that is, $dx$ and $dy$ are meromorphic but might have non-vanishing residues. Then the condition that the differential of $x^{\dagger}=x+\psi(y)$ is also meromorphic imposes certain conditions on $x$, $y$, and $\psi$.
In the presence of logarithmic singularities, it is more natural to consider not the traditional topological recursion but logarithmic topological recursion defined in Section 2.2, see also [ABDKS4]. The LogTR differentials comparing with TR differentials may posses a small number of extra poles with controlled principal parts. One of the main observations of [ABDKS4] is that the $x-y$ duality preserves LogTR rather than just TR.
Due to the fact that $y$ is multivalued when it has logarithmic singularities, it is tricky to formulate and prove a general theorem in the vein of Theorem 6.1 in this case, since we need to deal with the composition $\psi(y(z))$. So instead we are going to formulate a principle for which we aim in this case.
Principle 6.3.
Assume that $dx$ and $dy$ are meromorphic. Let $\psi$ be such that the composition $\psi(y(z))$ makes sense and $d\psi(y(z))$ is univalued and meromorphic. Let $x^{\dagger}=x+\psi(y)$ and assume that all zeros of $dx$, $dy$, and $dx^{\dagger}$ are simple, and that the zeros of $dx$ are disjoint from the zeros of $dy$, and $\psi$ is regular at the critical values of $y$.
Let $\tilde{\psi}$ be the $\hbar$-deformation of $\psi$ described below. Then the system of differentials $\{\omega^{(g)}_{n}\}$ solves LogTR for the spectral curve $(x,y)$ if and only if its symplectic dual system of differentials $\{\omega^{{\dagger},(g)}_{n}\}$ solves LogTR for the spectral curve $(x^{\dagger},y)$.
We do not prove the statement of this principle in general, but we prove it below for three large families of $(x,y,\psi)$ for which the conditions are satisfied.
6.2.1. Construction of $\tilde{\psi}$
The construction for $\tilde{\psi}$ goes as follows. Let $\{a_{1},\dots,a_{M}\}$ be the set of all simple poles of $d\psi(y(z))$ which are not simultaneously poles of $dy(z)$. Then $[\hbar^{2g}]\tilde{\psi}(y(z),\hbar)$, for $g\geq 1$, is the unique function (here $\tilde{\psi}(y(z),\hbar)$ is treated not as a composition but just as some function depending on $z$ and $\hbar$) such that $[\hbar^{2g}]\tilde{\psi}(y(z),\hbar)dy(z)$ does not have any poles apart from $a_{1},\dots,a_{M}$, and its principal part at $a_{i}$, for $i=1,\dots,M$, coincides with the principal part of
(83) $[\hbar^{2g}]\left(\frac{1}{\alpha_{i}\mathcal{S}(\alpha_{i}\hbar\partial_{y})}\log(z-a_{i})\right)dy,$
where $\alpha_{i}^{-1}$ is the residue of $d\psi(y(z))$ at $a_{i}$. It is not clear how to pass from $\tilde{\psi}(y(z),\hbar)$ to $\tilde{\psi}(\theta,\hbar)$ in general, but in the cases described below it is clear. In particular, if $M=0$ then $\tilde{\psi}(\theta,\hbar)=\psi(\theta)$.
6.2.2. Principle at work
We discuss below three cases (families) of $(x,y,\,\psi)$ satisfying the conditions of Principle 6.3. We do not claim that this list is exhaustive, but the respective theorems, in particular, imply that TR holds for (generating functions of) all generalized double Hurwitz numbers appearing in the literature, as we will see below. In all cases we assume that $\Sigma=\mathbb{C}\mathrm{P}^{1}$.
  1. Case I.
    Let $x$ be such that $dx$ is meromorphic, and
    (84) $y$ $=R(z),$
    (85) $\psi$ $=T(\theta)+\sum\limits_{i}\beta^{-1}_{i}\log(\theta-b_{i}),$
    where $R$ and $T$ are rational functions, and $\beta_{i}$ and $b_{i}$ are some constants, such that $y(z)$ and $\psi(\theta)$ are not constant functions. We also assume that for all $i$ expressions $R(z)-b_{i}$ have simple zeros and poles which are distinct over all $i$. We assume that $\psi(\theta)$ is regular at the critical values of $y$. Furthermore, we assume that all zeros of $dx$, $dy$, and $dx^{\dagger}$ are simple, and that the zeros of $dx$ are disjoint from the zeros of $dy$, then the zeros of $dx^{\dagger}$ are also disjoint from the zeros of $dy$.
  2. Case II.
    Let $x$ be such that $dx$ is meromorphic, and
    (86) $y$ $=R(z)+\sum\limits_{i}\alpha_{i}^{-1}\log(z-a_{i}),$
    (87) $\psi$ $=\varkappa_{1}\theta+\varkappa_{0},$
    where $R$ is a rational function and $\alpha_{i}$, $a_{i}$, and $\varkappa_{i}$ are some constants, such that $y(z)$ and $\psi(\theta)$ are not constant functions and $z=\infty$ is not a critical point of $y$. Furthermore, we assume that all zeros of $dx$, $dy$, and $dx^{\dagger}$ are simple, and that the zeros of $dx$ are disjoint from the zeros of $dy$, then the zeros of $dx^{\dagger}$ are also disjoint from the zeros of $dy$.
  3. Case III.
    Let $x$ be such that $dx$ is meromorphic, and
    (88) $y$ $=\sum\limits_{i}\alpha_{i}^{-1}\log(z-a_{i}),$
    (89) $\psi$ $=T(e^{\gamma_{1}\theta},\dots,e^{\gamma_{K}\theta})+\varkappa_{1}\theta+\varkappa_{0},$
    where $T$ is a rational function in $K$ variables and $\kappa_{i}$, $\alpha_{i}$ and $\gamma_{i}$ are some constants such that $\forall i,j\;\,{\dfrac{\gamma_{i}}{\alpha_{j}}\in\mathbb{Z}}$, and $y(z)$ and $\psi(\theta)$ are not constant functions. Furthermore, we assume that $\psi$ is regular at the critical values of $y$, that all zeros of $dx$, $dy$, and $dx^{\dagger}$ are simple, and that the zeros of $dx$ are disjoint from the zeros of $dy$, then the zeros of $dx^{\dagger}$ are also disjoint from the zeros of $dy$.
Remark 6.4.
The restriction on zeros and poles of $R-b_{i}$ in Case I. comes from the fact that if we do not impose it, then the procedure for determining $\tilde{\psi}$ described above gives us $\tilde{\psi}(y(z))$ but then we cannot represent it as a composition of $\tilde{\psi}(\theta)$ and $y(z)$ for some good $\tilde{\psi}(\theta)$.
6.3. Case I
Following the recipe described after Principle 6.3, for Case I. we obtain
(90) $\tilde{\psi}=T(\theta)+\sum_{i}\frac{1}{\beta_{i}\mathcal{S}(\beta_{i}\hbar\partial_{\theta})}\log(\theta-b_{i}).$
Theorem 6.5.
Let $(\Sigma,\,x,\,y,\,\psi)$ be as in Case I. Let $\tilde{\psi}$ is given by (90). Then the system of differentials $\{\omega^{(g)}_{n}\}$ solves TR for the spectral curve $(x,y)$ if and only if its symplectic dual system of differentials $\{\omega^{{\dagger},(g)}_{n}\}$ solves TR for the spectral curve $(x^{\dagger},y)$.
Proof.
We follow the lines of the proof of Theorem 6.1, and the notation follows the notation of Theorem 5.1 (in particular, see (66) and (67)).
Transformation (a) of Theorem 5.1 preserves LogTR by Theorem 3.10 (part 7.). Let us see that after the application of transformation (b) of Theorem 5.1 LogTR still holds.
Indeed, the linear loop equations for $(g,n)=(0,1)$ still hold due the fact that $\psi(\theta)$ is regular at the critical values of $x^{\vee}=x^{\dagger\vee}$ and thus $\psi(x^{\vee}(z))dx^{\vee}$ is regular and vanishing at the critical points of $x^{\vee}=x^{\dagger\vee}$. For $g>1$ we have $\omega^{{\dagger}\vee,(g)}_{1}-\omega^{\vee,(g)}_{1}=\Bigl{(}[\hbar^{2g}]\tilde{\psi}(x^{\vee}(z),\hbar)\Bigr{)}\,dx^{\vee},$ where $[\hbar^{2g}]\tilde{\psi}(x^{\vee}(z),\hbar)\sim(x^{\vee}(z)-b_{i})^{-2g}$ which is regular at the critical points of $x^{\vee}=x^{\dagger\vee}$ since $b_{i}$ do not coincide with critical values of $x^{\vee}=x^{\dagger\vee}$; thus linear loop equations still hold. Regarding the quadratic loop equations, the same reasoning as in Theorem 6.1 works, as $\Bigl{(}[\hbar^{2g}]\tilde{\psi}(x^{\vee}(z),\hbar)\Bigr{)}\,dx^{\vee}$ is symmetric under the deck transformations and thus gets killed by $\Delta$.
The projection property is not completely trivial now as not only the $\omega^{(0)}_{1}$ differential gets changed, but $\omega^{(g)}_{1}$ for $g>0$ as well. The only poles that $\omega^{{\dagger}\vee,(g)}_{1}-\omega^{\vee,(g)}_{1}=\Bigl{(}[\hbar^{2g}]\tilde{\psi}(x^{\vee}(z),\hbar)\Bigr{)}\,dx^{\vee}$ has are at the zeros and poles of the expressions $R(z)-b_{i}$, and (90) guarantees that the principal parts there agree with (5), where $\alpha_{i}^{-1}=\beta_{i}^{-1}$ is the residue of $\psi(x^{\vee}(z))$ at the given pole, due to the assumption that the zeros and poles of $R(z)-b_{i}$ are simple and distinct. Thus the logarithmic projection property holds.
Thus, after transformation (b) LogTR still holds, and after transformation (c) it holds as well again due to Theorem 3.10 (part 7.), where we again note that the zeros of $dx^{\dagger}$ are disjoint from the zeros of $dy$, same as in Theorem 5.1. ∎
Corollary 6.6 (Extended Family I of [BDKS2]).
In the case of $x(z)=\log z$ we obtain a new proof of topological recursion for the (extended version of) Family I of Orlov–Scherbin hypergeometric tau functions from [BDKS2].
Proof.
Follows from Theorem 6.5 and Proposition 4.4. Family I of [BDKS2] itself corresponds to setting $\beta_{i}=\pm 1$ and requiring that $T$ is a polynomial and that $y(0)=0$ and $\psi(0)=0$. ∎
6.4. Case II
Theorem 6.7.
Let $(\Sigma,\,x,\,y,\,\psi)$ be as in Case II. Let $\tilde{\psi}=\psi$. Then the system of differentials $\{\omega^{(g)}_{n}\}$ solves LogTR for the spectral curve $(x,y)$ if and only if its symplectic dual system of differentials $\{\omega^{{\dagger},(g)}_{n}\}$ solves LogTR for the spectral curve $(x^{\dagger},y)$.
Proof.
Note that all new logarithmic singularities in $y^{\dagger\vee}$ in this case are not vital (they coincide with the singularities of $x^{\dagger\vee}$), so the proof is completely analogous to the proof of Theorem 6.1. ∎
Corollary 6.8 (Extended Family II of [BDKS2]).
In the case of $x(z)=\log z$ we obtain a new proof of topological recursion for the Family II of Orlov–Scherbin hypergeometric tau functions from [BDKS2].
Proof.
Follows from Theorem 6.7 and Proposition 4.4. Family II of [BDKS2] itself corresponds to setting $\alpha_{i}=\pm 1$, $\varkappa_{0}=0$, and requiring that $y(0)=0$ and $y$ does not have a vital singularity at $\infty$ with respect to $x^{\dagger}$. ∎
The topological recursion for an extension of Family II of [BDKS2] was also earlier obtained in [kramer2022kp].
6.5. Case III
Theorem 6.9.
Let $(\Sigma,\,x,\,y,\,\psi)$ be as in Case III. Let $\tilde{\psi}=\psi$. Then the system of differentials $\{\omega^{(g)}_{n}\}$ solves LogTR for the spectral curve $(x,y)$ if and only if its symplectic dual system of differentials $\{\omega^{{\dagger},(g)}_{n}\}$ solves LogTR for the spectral curve $(x^{\dagger},y)$.
Proof.
Note that all new logarithmic singularities in $y^{\dagger\vee}$ in this case are not vital (they coincide with the singularities of $x^{\dagger\vee}$), so the proof is completely analogous to the proof of Theorem 6.1. ∎
Corollary 6.10 (Extended Family III of [ABDKS4, Section 3.11]).
In the case of $x(z)=\log z$ we obtain a new proof of topological recursion for the Family III of Orlov–Scherbin hypergeometric tau functions from [ABDKS4, Section 3.11].
Proof.
Follows from Theorem 6.9 and Proposition 4.4. Family III of [ABDKS4, Section 3.11] itself corresponds to setting $\alpha_{i}=\pm 1$, $\kappa_{i}=0$ and $T(\zeta)=\alpha(\zeta-1)$, and requiring that $y(0)=0$. ∎
7. More examples
In this section we present several examples supplementing those in [ABDKS4] and demonstrating interaction between the notions of $x-y$ duality, symplectic duality, topological recursion, and LogTR.
7.1. Topological recursion with unramified $y$-function
Proposition 3.1 of [ABDKS4] provides explicit formulas for TR differentials for the spectral curves of genus zero and $(x,y)$ with arbitrary $x$ such that $dx$ is rational and with unramified $y$, that is, $dy$ never vanishes and $y$ can be parametrized either as $y=z$ or as $y=\log z$ with a global coordinate $z$. These formulas express these differentials as $x-y$ dual to the trivial LogTR differentials for the corresponding spectral curves $(y,x)$. Namely, we have
(91) $y$ $=z:$ $\omega^{(g)}_{n}$ $=[\hbar^{2g-2+n}](-1)^{n}\prod_{i=1}^{n}\Bigl{(}\sum_{r=0}^{\infty}\bigl{(}-d_{i}\tfrac{1}{dx_{i}}\bigr{)}^{r}[u_{i}^{r}]e^{u_{i}(\mathcal{S}(u_{i}\hbar\partial_{z_{i}})\hat{x}_{i}-x_{i})}dz_{i}\Bigr{)}$
$\qquad\qquad(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{z_{i}+\frac{u_{i}\hbar}{2}-z_{\sigma(i)}+\frac{u_{\sigma(i)}\hbar}{2}},$
(92) $y$ $=\log z:$ $\omega^{(g)}_{n}$ $=[\hbar^{2g-2+n}](-1)^{n}\prod_{i=1}^{n}\Bigl{(}\sum_{r=0}^{\infty}\bigl{(}-d_{i}\tfrac{1}{dx_{i}}\bigr{)}^{r}[u_{i}^{r}]e^{u_{i}(\mathcal{S}(u_{i}\hbar z_{i}\partial_{z_{i}})\hat{x}_{i}-x_{i})}dz_{i}\Bigr{)}$
$\qquad\qquad(-1)^{n-1}\!\!\sum_{\sigma\in\{n\text{-cycles}\}}\prod_{i=1}^{n}\frac{1}{e^{\frac{u_{i}\hbar}{2}}z_{i}-e^{-\frac{u_{\sigma(i)}\hbar}{2}}z_{\sigma(i)}},$
here $\hat{x}(z,\hbar)=\sum_{i=1}^{M}\frac{1}{\mathcal{S}(\alpha_{i}\hbar\partial_{y_{i}})}\frac{\log(z-a_{i})}{\alpha_{i}}+R(z)$, $a_{1},\dots,a_{M}$ are the LogTR-vital singularities of $x$ and $1/\alpha_{i}$ are the corresponding residues of $dx$. The term $R(z)$ may contain contributions of other non-LogTR-vital logarithmic singularities, so we do not assume that $R$ is rational.
To be precise, Eq. (91) and (92) provide the LogTR differentials which coincide, in fact, with TR differentials in the case $y=z$ for any $x$ and also in the case $y=\log z$ if $dx$ has a pole both at $z=0$ and $z=\infty$. The next theorem shows that these formulas are contained implicitly in the closed expression for the Orlov–Scherbin family derived earlier in [BDKS2] and (51).
Proposition 7.1.
1. If $y=z$, then the differentials (91) can be obtained as a special case of the Orlov–Scherbin differentials (51) with
(93) $\hat{y}(z)=z,\qquad\tilde{\psi}(\theta)=\frac{1}{\mathcal{S}(\hbar\partial_{\theta})}\log\theta-\hat{x}(\theta).$
2. If $y=\log z$, then the differentials (92) can be obtained as a special case of the Orlov–Scherbin differentials (51) with
(94) $\hat{y}(z)=\log z-\hat{x}(z),\qquad\tilde{\psi}(\theta)=\theta.$
In both cases Eq. (51) should be applied formally ignoring the condition $y(0)=\psi(0)=0$ required for the Orlov–Scherbin family. It follows that this formula provides the correct answer even though the corresponding differentials could be singular at $z=0$ and have no VEV presentation.
Proof.
In the case $y=z$ we treat expressions for the differentials of the Orlov–Scherbin family as a special case of symplectic duality and apply Theorem 5.1. We start with the trivial differentials associated with the spectral curve $(z,\log z)$. On the first step we apply $x-y$ duality and obtain LogTR differentials for the spectral curve $(\log z,z)$ that we denote by $\tilde{\omega}^{(g)}_{n}$. By Example 3.11 we know that they are trivial for $n>1$ and in the case $n=1$ they are given by (34),
(95) $\tilde{\omega}^{(g)}_{1}=\Bigl{(}[\hbar^{2g}]\tfrac{1}{\mathcal{S}(\hbar\partial_{z})}\log z\Bigr{)}\;dz.$
On the next step we modify these $(g,1)$ differentials by adding terms corresponding to the chosen function $\tilde{\psi}$ and obtain the differentials $\omega^{\vee,(g)}_{n}$ which are trivial for $n>1$ and in the case $n=1$ given by
(96) $\omega^{\vee,(g)}_{1}=\tilde{\omega}^{(g)}_{1}-[\hbar^{2g}]\tilde{\psi}(z)\;dz=[\hbar^{2g}]\hat{x}\;dz.$
These are exactly the differentials $x-y$ dual to the differentials (91). Therefore, when we apply $x-y$ duality to these differentials we obtain exactly the differentials (91).
In the case $y=\log z$ the argument is even simpler. We notice that in the case $\tilde{\psi}(\theta)=\theta$ we have
(97) $e^{-v\psi(\theta)}\partial_{\theta}^{r}e^{v\mathcal{S}(v\hbar\partial_{\theta})\tilde{\psi}(\theta)}=v^{r},\qquad(\mathcal{S}(u_{i}\hbar z\partial_{z})-1)\log z=0.$
Therefore, the right hand side of (51) simplifies to (92) with $u_{i}$ replaced by $-u_{i}$. ∎
7.2. Shifted Kontsevich–Witten potential
The Kontsevich–Witten (KW) partition function can be computed by TR with the spectral curve $(x,y)=(z^{2}/2,z)$. We would like to relate this TR to the TR differentials of the Orlov–Scherbin family. One of the ways to do that is to apply Proposition 7.1 which claims that the TR differential for the spectral curve $(z^{2}/2,z)$ can be computed by the closed formulas (51) for the Orlov–Scherbin family with the following specialization of parameters:
(98) $\hat{y}(z)=y(z)=z,\qquad\tilde{\psi}(\theta)=\frac{1}{\mathcal{S}(\hbar\partial_{\theta})}\log\theta-\frac{\theta^{2}}{2}.$
This specialization does not satisfy the requirement $\psi(0)=0$ and does not imply the Orlov–Scherbin type of the associated tau function. Another way to relate the KW partition function to the Orlov–Scherbin family is to modify slightly the function $y$, which corresponds to a certain shift of times in the KW partition function. For example, consider the following spectral curve data
(99) $\left(\frac{z^{2}}{2},~{}\frac{z}{1-a^{2}z^{2}}\right),$
where $a\neq 0$ is a complex parameter. The initial TR differential for this spectral curve is
(100) $\omega^{(0)}_{1}=\frac{z}{1-a^{2}z^{2}}\,d\Bigl{(}\frac{z^{2}}{2}\Bigr{)}=z\;d\bigl{(}-\tfrac{1}{2a^{2}}\log(1-a^{2}z^{2})\bigr{)}$
and, since the zeros of the differentials $d(z^{2}/2)=zdz$ and $d\bigl{(}-\tfrac{1}{2a^{2}}\log(1-a^{2}z^{2})\bigr{)}=zdz/(1-a^{2}z^{2})$ coincide, it shows that the TR for the spectral curves $(z^{2}/2,z)$ and $\bigl{(}-\tfrac{1}{2a^{2}}\log(1-a^{2}z^{2}),\;z\bigr{)}$ also coincide. This leads to the following variety of possibilities for computing these differentials, see [BM, ZhouKW, Eynard-intersections, Chen, DKPS-cut-and-join, AleKPIII, KN, ABDKS4].
Proposition 7.2.
The following five ways for the computation of the TR differentials $\omega^{(g)}_{n}$ associated with the spectral curve (99) are equivalent to one another:
  1. (1)
    apply directly the TR procedure for the spectral curve data (99) in a straightforward way;
  2. (2)
    use the following explicit expressions for these differentials in terms of intersection numbers of $\psi$ and $\kappa$ classes over moduli spaces:
    (101) $\xi_{k}(z)$ $=\Bigl{(}-\frac{1}{z}\frac{d}{dz}\Bigr{)}^{k}\frac{1}{z}=\frac{(2k-1)!!}{z^{2k+1}},$
    (102) $\omega^{(g)}_{n}$ $=\sum_{k_{1},\dots,k_{n}}\int_{\overline{\mathcal{M}}_{g,n}}e^{\sum_{k=1}^{\infty}s_{k}a^{2k}\kappa_{k}}\psi_{1}^{k_{1}}\dots\psi_{n}^{k_{n}}\prod_{i=1}^{n}d\xi_{k_{i}}\!(z_{i}),$
    where the constants $s_{k}$ are determined by the expansion $e^{-\sum_{k=1}^{\infty}s_{k}t^{k}}=\sum_{k=0}^{\infty}(2k+1)!!t^{k}$;
  3. (3)
    apply $x-y$ duality formula (91) to the spectral curve
    (103) $(x,y)=\bigl{(}-\tfrac{1}{2a^{2}}\log(1-a^{2}z^{2}),\;z\bigr{)}$
    with
    (104) $\hat{x}(z)=\frac{1}{\mathcal{S}(2a^{2}\hbar\partial_{z})}x(z);$
  4. (4)
    use an expression for these differentials in terms of triple Hodge integrals associated with the spectral curve of item (3):
    (105) $\xi_{k}(z)$ $=\Bigl{(}-\frac{d}{dx(z)}\Bigr{)}^{k}\frac{1}{z},$
    (106) $\Lambda(q)$ $=\sum_{k=1}^{g}q^{k}\lambda_{k},$
    (107) $\omega^{(g)}_{n}$ $=\sum_{k_{1},\dots,k_{n}}\int_{\overline{\mathcal{M}}_{g,n}}\Lambda(-2a^{2})\Lambda(-2a^{2})\Lambda(a^{2})e^{\sum_{k=1}^{\infty}\frac{(-1)^{k}a^{2k}}{k}\kappa_{k}}\psi_{1}^{k_{1}}\dots\psi_{n}^{k_{n}}\;\prod_{i=1}^{n}d\xi_{k_{i}}\!(z_{i});$
  5. (5)
    apply Eq. (51) for the Orlov–Scherbin family with the following specialization of parameters:
    (108) $\hat{y}(\zeta)=y(\zeta)=\zeta,\qquad\psi(\theta)=-\log(1-a^{3}\theta),\quad\tilde{\psi}(\theta)=\frac{1}{\mathcal{S}(\hbar\partial_{\theta})}\psi(\theta),$
    (109) $X=\zeta\,e^{-\psi(y(\zeta))}=\zeta-a^{3}\zeta^{2},$
    where $\zeta=\frac{1-a\,z}{2a^{3}}$ is a new affine coordinate on $\Sigma=\mathbb{C}P^{1}$.
The equivalence of all these relations is known, we just collected them together in a compact form. The parameters $(-2a^{2},-2a^{2},a^{2})$ of triple Hodge integrals are inverse to the residues $\bigl{(}\frac{-1}{2a^{2}},\frac{-1}{2a^{2}},\frac{1}{a^{2}}\bigr{)}$ of the differential $dx$ of item (3) at its three poles. The triple Hodge integrals without insertion of $\kappa$ classes correspond to the choice of $y$ such that $\frac{dy}{dx}=\frac{1}{z}$, that is, $y=\frac{1}{2a}\log\frac{1+az}{1-az}$. Since our choice $y=z$ is different, an extra insertion of $\kappa$ classes is needed.
Note also that the computation of item (5) shows that the coefficients of the power expansion of $\omega^{(g)}_{n}$ at the point $z_{i}=1/a$ in the local coordinates $\zeta_{i}$ are monotone Hurwitz numbers.
7.3. Atlantes vs spin Hurwitz numbers
Consider the differentials of the Orlov–Scherbin family with
(110) $\hat{y}=y=z.$
If $e^{\psi(\theta)}$ is rational (and such that all its zeros and poles distinct from 00 and $\infty$ are simple) then (51) with
(111) $X=\frac{z}{e^{\psi(y(z))}},\qquad\tilde{\psi}(\theta)=\frac{1}{\mathcal{S}(\hbar\partial_{\theta})}\psi(\theta)$
produces TR differentials for the spectral curve
(112) $x=\log X=\log z-\psi(z),\quad y(z)=z,$
see [ACEH, BDKS2]. By Proposition 7.1, these differentials can also be given by (91) with
(113) $\hat{x}=\frac{1}{\mathcal{S}(\hbar\partial_{z})}(\log z-\psi(z)).$
Assume now that $\psi(\theta)$ is itself rational, for instance, $\psi(\theta)=\theta^{r}$. Then the differentials produced by the same formula are not necessarily TR differentials any more: they may attain extra poles at $z=\infty$. The numbers appearing in the power expansion of these differentials at the origin in the local coordinates $X_{i}$ are known as Atlantes Hurwitz numbers [ALS]. In order to correct the situation and to obtain the true TR differentials for the same spectral curve, one should apply (51) to a different $\hbar$-deformation of $\psi$, namely, to $\tilde{\psi}(\theta)=\psi(\theta)$ (we still assume that $\psi(\theta)$ is rational). Equivalently, we can apply (91) with
(114) $\hat{x}=\frac{1}{\mathcal{S}(\hbar\partial_{z})}\log z-\psi(z).$
The corresponding differentials enumerate spin Hurwitz numbers.
7.3.1. Transalgebraic topological recursion
In order to reproduce the Atlantes Hurwitz numbers in the case when $\psi(\theta)=\theta^{r}$, it was suggested in [BKW] to consider the following family of spectral curves
(115) $X_{N}(z)$ $=z\Big{(}1+\frac{\psi(z)}{N}\Big{)}^{-N};$ $Y_{N}(z)=\frac{z}{X_{N}(z)};$ $(\omega_{N})^{(0)}_{1}=Y_{N}dX_{N},$
depending on a parameter $N$; compute the differentials $(\omega_{N})^{(g)}_{n}$ and take the limit $N\to\infty$ (it is a special case of the proposal made in [BKW] under the name transalgebraic topological recursion). The proof that the limit procedure works in this case (and gives the desired differentials) is given in [BKW], and it is a quite subtle analytic argument. The methods of the present paper give an alternative proof of their main theorem.
7.3.2. Atlantes Hurwitz numbers
In order to apply our technique to the limit procedure suggested in [BKW], we have to pass from $X_{N}$ and $Y_{N}$ to $x_{N}=\log X_{N}$ and $y_{N}=z$. If we do it naively, we loose some of the critical points of $dX_{N}$ (the situation is similar to the one for Bousquet-Mélou–Schaeffer differentials discussed in [BDS, ABDKS4]). To remedy this problem, we assume that $N$ is a positive integer and introduce an $\epsilon$-deformation of the family of differentials constructed in [BKW] by
(116) $\psi_{N,\epsilon}(z)$ $=\sum\nolimits_{i=1}^{N}\,\log P_{i,\epsilon}(z),$
(117) $x_{N,\epsilon}(z)$ $=\log z-\psi_{N,\epsilon}(z)=\log z-\sum\nolimits_{i=1}^{N}\,\log P_{\epsilon,i}(z),$
(118) $y_{N,\epsilon}(z)$ $=z,$
where $P_{1,\epsilon}\dots,P_{N,\epsilon}$ are small independent polynomial deformations of the polynomial $1+N^{-1}\psi(z)$ depending on a small parameter $\epsilon$, such that all zeros of $P_{1,\epsilon}\dots,P_{N,\epsilon}$ should be simple and close to the zeros of $1+N^{-1}\psi(z)$. The differential $dx_{N,\epsilon}$ is meromorphic and the topological recursion is well defined. We obviously have
(119) $\lim\limits_{N\to\infty}\lim\limits_{\epsilon\to 0}\psi_{N,\epsilon}(\theta)=\psi(\theta),\qquad\lim\limits_{N\to\infty}\lim\limits_{\epsilon\to 0}dx_{N,\epsilon}(z)=dx(z).$
According to [ABDKS4, Proposition 3.1], the TR differentials for this family of spectral curves are given by (91) with the function $\hat{x}_{N,\epsilon}$ taken in the form
(120) $\hat{x}_{N,\epsilon}(z)=\frac{1}{\mathcal{S}(\hbar\partial_{z})}x_{N,\epsilon}(z).$
The formula is manifestly compatible with the limits $\epsilon\to 0$ and $N\to\infty$, where the first limit gives the $(\omega_{N})^{(g)}_{n}$ differentials of [BKW] (it is a case of the standard compatibility of topological recursion with the limits discussed in detail in [limits]), and the subsequent $N\to\infty$ gives the formulas for the differentials computed with $\hat{x}$ given by (113), that is, the Orlov–Scherbin differentials enumerating Atlantes Hurwitz numbers (and not satisfying the topological recursion).
7.3.3. $r$-spin Hurwitz numbers
Let us also illustrate how one can obtain the differentials enumerating $r$-spin Hurwitz numbers via a similar limit procedure. In this case the computation below is not reproducing the transalgebraic topological recursion of [BKW] as we loose the contribution of some of the critical points that they have for finite $N$.
Let $\psi(\theta)=\theta^{r}$ and consider
(121) $\psi_{N}(\theta)$ $=N\log(1+N^{-1}\psi(\theta)),$
(122) $x_{N}(z)$ $=\log z-\psi_{N}(z),$
(123) $y_{N}(z)$ $=z.$
Then, $dx_{N}=\frac{dz}{z}-\frac{\psi^{\prime}(z)\,dz}{1+N^{-1}\psi(z)}$ is rational and we have obviously
(124) $\lim\limits_{N\to\infty}\psi_{N}(\theta)=\psi(\theta),\qquad\lim\limits_{N\to\infty}dx_{N}(z)=dx(z).$
It is important to stress that the set of zeros of $dx_{N}$ does differ from the set of zeros of $dX_{N}$ given in (115), so the TR differentials for this family differ from the ones considered in the transalgebraic TR approach.
According to [ABDKS4, Proposition 3.1], the TR differentials for this family of spectral curves are given again by (91) with the function $\hat{x}_{N}$ taken in the form
(125) $\hat{x}_{N}(z)=\frac{1}{\mathcal{S}(\hbar\partial_{z})}\log z-\frac{1}{\mathcal{S}(N^{-1}\hbar\partial_{z})}\psi_{N}(z).$
The differentials computed by this formula have a limit as $N\to\infty$, and the limiting differentials are given also by (91) with the function $\hat{x}$ given by (114), that is, these are differentials for $r$-spin Hurwitz numbers.
[…]